### v1

# Rethinking Weight-Averaged Model-merging

Abstract
Weight-averaged model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without fine-tuning or retraining. However, the underlying mechanisms that explain its effectiveness remain largely unexplored. In this paper, we investigate this technique from three novel perspectives to provide deeper insights into how and why weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns; (2) we investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales. Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process.
1 Introduction
Model-merging can combine multiple independently trained models into a single, efficient model without introducing extra inference computations. This approach has been applied in contexts such as natural language processing [36], computer vision [14], and their sub-areas (e.g. federated learning [30], distillation [11], adversarial learning [42, 3] and large language model [20, 43]), where the ability to merge models effectively can lead to improved model performance, efficiency, and flexibility in deployment.
Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective, especially weight-averaged model-merging [33] (e.g. Uniform Soups and Greedy Soups) as the mainstream techniques. Without a clear understanding of the underlying mechanisms of model-merging, researchers and practitioners may lack guidance on when and how to apply it in a principled way. This research gap motivates the need to systematically explore the underlying patterns and behaviors in weight-averaged model-merging.
In this paper, we address this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
  • An examination of the patterns contained in the averaged model weights through the visualization of the patterns present in both linear layer as classifiers or inside deep learning models used on multiple datasets. Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
  • An investigation of averaging on weights versus averaging on features strategies allows us to evaluate the relative effectiveness and limitations of model-merging and traditional model ensembles across multiple architectures and datasets, uncovering distinct benefits and behaviors;
  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of of model regularization.
Based on our viewpoints and findings, we conclude with useful proposals and source code that will be publicly available to provide a good reference for the model-merging research community.
2 Related Work
(a) Average of all images for each class.
(b) Linear classifiers visualization.
Figure 1: Average of all images for each class and linear classifiers visualization for each class in Cifar100 [12] dataset. The detailed experimental analysis is in Sec. 6.1.
Figure 2: Merge the linear classifiers for the first 5 classes (within the above 10 classes) with the last 5 classes on Cifar100. It interprets the reason why the model-merging techniques can work as they essentially operate the templates contained in model weights.
In recent years, model-merging [26, 17, 21, 4, 2, 19, 23, 35] has attracted growing interest. This approach has led to a range of methods focused on averaging or merging weights from independently trained models to enhance model performance or generalization. For example, Wortsman et al. [33] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup”. Further, Yadav et al. proposed Ties-Merging [34] based on parameter averaging by considering conflicts between the model weights during training, allowing for smoother parameter alignment across models for merging. Some other approaches [1, 25] address the problem from a different perspective, focusing on finding weight merging operations via the similarity analyses of model features. Yu et al. [39] combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper. To focus on a specific scope, we aim to offer a more thorough analysis of weight-averaged model-merging than the study in [33], which serves as the foundation work for the area.
3 The Patterns Contained in Model Weights
In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching. Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with inner product which is a widely explored operation in neural networks. By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
3.1 Connections Between Inner Product and Template Matching
The weights of the model can be interpreted as templates or prototypes for the respective class activation. A couple of decades ago, Adaboost [7] adopted heuristically designed patterns such as HAAR features [29], which are simple, hand-crafted filters that capture edges, lines, and other low-level image characteristics through weighted sums and differences in pixel intensities. During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in the Fig. 6.1. A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data. When an image is presented to the model, its activation is determined by the inner product between the image’s feature representation and each class’s template. The class whose template yields the highest activation is selected as the predicted class, indicating that the template matches the input image most closely. This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different “important” features that contribute the most towards the objectives.
This inner product computation is essentially a similarity measurement. In this context, the averaging operation in the weight space represents a linear combination of weight matrices (shown as Fig. 2), which can be seen as Mixup [40] operation in the weight space. The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space. This also explains why recent research [5, 41, 10] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
This concept extends to model-merging, where multiple models or variations of a model are combined. In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model. The model after merging can work more effectively than individuals because they represent a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
3.2 Formal Representation from the Perspective of Template Matching
For a linear classifier, each row of the weight vector $\mathbf{w}\in\mathbb{R}^{n}$ can be interpreted as a template for a particular class, and the input vector is denoted as $\mathbf{x}\in\mathbb{R}^{n}$. The inner product between $\mathbf{w}$ and $\mathbf{x}$ measures the alignment between them:
$\small\mathbf{w}\cdot\mathbf{x}=\sum_{i=1}^{n}w_{i}x_{i}=\|\mathbf{w}\|\|\mathbf{x}\|\cos\theta,$ (1)
where $\|\mathbf{w}\|$ and $\|\mathbf{x}\|$ are the norms of $\mathbf{w}$ and $\mathbf{x}$, and $\theta$ is the angle between the two vectors. When $\theta=0$, then $\cos\theta=1$, which means that we have $\mathbf{w}\cdot\mathbf{x}=\|\mathbf{w}\|\|\mathbf{x}\|$, resulting in a weight matrix $\mathbf{w}$ that perfectly matches the input $\mathbf{x}$, as the inner product in (1) is maximized when the template $\mathbf{w}$ is perfectly aligned with the input $\mathbf{x}$. On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned. This mechanism also applies for the linear layers within deep learning models.
{mdframed}
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 1: From the analysis above and the experimental study shown in Sec. 6.1, since weight matrices contain generalizable template information, if the merged model’s performance is suboptimal, more focus should be directed toward resolving conflicts within the individual learned patterns in the weights.
4 Averaging on Weights vs. Averaging on Features
Figure 3: The 10 model-merging/ensemble accuracy comparison between different architectures on Cifar100. Comparison between averaging on Weights (Uniform Soups, Greedy Soups), averaging on features (Ens Logits, Ens Features, Greedy Ens Logits, Greedy Ens Features) and original model performance (Performance Average) on Cifar100 using many different deep learning architectures, including CNNs and ViTs. The detailed experimental analysis is in Sec. 6.2.
Averaging on weights (model-merging) and Averaging on features (model ensembling) are two different ways to aggregate knowledge from multiple models without retraining. These methods have distinct properties and lead to different testing behaviors. In this section, we explore the differences between these two approaches and analyze their effects on the model effectiveness and efficiency, one example experiment is shown in Fig. 3.
4.1 Linear Model Scenarios
Averaging on Weights. Assuming that our model only contains linear layers, averaging on weights can be referred to as a process where, given two weight matrices $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, we compute their average first, and then apply the result to the input feature vector $\mathbf{x}$:
$\small\overline{\mathbf{W}}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)$ (2)
The resulting output feature vector is then computed as:
$\small\overline{\mathbf{h}}=\overline{\mathbf{W}}\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}$ (3)
Since both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ represent linear transformations, the overall transformation remains linear. When no activation function is involved, this method is mathematically equivalent to averaging on features, as discussed in the next subsection.
Averaging on features involves first applying each weight matrix $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ to the input $\mathbf{x}$ separately, with:
$\small\mathbf{h}_{1}=\mathbf{W}_{1}\mathbf{x},\quad\mathbf{h}_{2}=\mathbf{W}_{2}\mathbf{x},$ (4)
and then averaging the resulting outputs:
$\small\overline{\mathbf{h}}^{\prime}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x}\right).$ (5)
In the context of linear operations, the methods in (3) and (5) are equivalent and yield the same result $\overline{\mathbf{h}}=\overline{\mathbf{h}}^{\prime}$
4.2 Impact of Non-linear Activation Functions
The equivalence between averaging on weights or on features breaks when non-linear activation functions $\phi(\cdot)$, such as ReLU or Sigmoid, are introduced. In the case of averaging on weights, the non-linear activation is applied after averaging the weight matrices:
$\small\overline{\mathbf{h}}=\phi\left(\overline{\mathbf{W}}\mathbf{x}\right)=\phi\left[\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}\right].$ (6)
Here, the non-linearity is applied to the weighted average of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. When averaging on features, the activation function is applied to each output feature independently:
$\small\mathbf{h}_{1}=\phi\left(\mathbf{W}_{1}\mathbf{x}\right),\text{ and }\mathbf{h}_{2}=\phi\left(\mathbf{W}_{2}\mathbf{x}\right).$ (7)
The final output is then the average of these activated features:
$\small\overline{\mathbf{h}}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left[\phi\left(\mathbf{W}_{1}\mathbf{x}\right)+\phi\left(\mathbf{W}_{2}\mathbf{x}\right)\right].$ (8)
Hence, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models will no longer hold for non-linear models. If we take ReLU for instance, each weight matrix retains its positive contribution through non-linear activation functions, while squashing to zero the negative contributions. As a result, averaging on features tends to preserve more information than averaging on weights, especially in scenarios where the transformations $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit diverse activations, such as when $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ have different signs, where their average may tend towards zero, leading to potential loss of information after the non-linear activation.
The key difference between the two approaches lies in how they interact with non-linear activation functions. Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect. However, this smoothing may also lead to a loss of fine-grained feature details, especially after non-linear activation, as illustrated in Eq. 6. In contrast, averaging on features allows each weight matrix to independently influence the final output after activation, also better adapting to individual data samples, preserving more variability and potentially enhancing the model’s expressive capacity. However, model ensemble will increase the computational cost linearly with the number of ensembled models; while model-merging will not. The FLOPs of models considered in the paper can be found in the supplementary materials.
{mdframed}
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analysis above and the experimental study shown in Sec. 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights. However, the averaging on weights can show relatively good performance without introducing extra inference computational cost. Hence, this forms a trade-off between computation usage and model performance when deciding between model-merging or model ensemble.
5 The Model Predictions and Weight Magnitudes
Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes on Cifar100. The values on the X and Y axis represent the accuracies. The experimental analysis is in Sec. 6.3.
In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how the model-merging works as a similar way of regularization.
5.1 Analysis of Weight Magnitude After Averaging
In this section, we analyze the effect of averaging the weights of two independent models, denoted as $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$, on their respective magnitudes. We define the maximum magnitude of the weights from both models as $M=\max(|\mathbf{W}_{A}|,|\mathbf{W}_{B}|)$. When we compute the averaged weights, we obtain:
$\small\overline{\mathbf{W}}=\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}.$ (9)
We seek to understand the relationship between the magnitude of the averaged weights $|\overline{\mathbf{W}}|$ and the maximum magnitude $M$.
Mathematical Justification. Using the triangle inequality [28], we establish:
$\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right\|\leq\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2},$ (10)
[…]
$\small\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2}\leq\max(\|\mathbf{W}_{A}\|,\|\mathbf{W}_{B}\|)=M.$ (11)
Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$. This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound equal to the maximum magnitude of the original weights, but in most cases it will be less than that maximum. Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
5.2 Variance Reduction in Model Weight Averaging
Consider two models, each with corresponding weight vectors $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. Let us denote the means and variances of the weight distributions for $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ as $\mathbb{E}[\mathbf{W}_{A}]=\mu_{A}$, $\text{Var}(\mathbf{W}_{A})=\sigma_{A}^{2}$ and $\mathbb{E}[\mathbf{W}_{B}]=\mu_{B}$, $\text{Var}(\mathbf{W}_{B})=\sigma_{B}^{2}$. The average of the weights of models A and B is defined as $\overline{\mathbf{W}}=(\mathbf{W}_{A}+\mathbf{W}_{B})/2$
Our goal is to compute the variance of $\overline{\mathbf{W}}$ and compare it to the variances of $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The expectation of the averaged weights $\overline{\mathbf{W}}$ is given by:
$\small\mathbb{E}[\overline{\mathbf{W}}]=\mathbb{E}\left[\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right]=\frac{\mathbb{E}[\mathbf{W}_{A}]+\mathbb{E}[\mathbf{W}_{B}]}{2}=\frac{\mu_{A}+\mu_{B}}{2}$ (12)
For the variance of $\overline{\mathbf{W}}$, we have:
$\small\text{Var}(\overline{\mathbf{W}})=\text{Var}\left(\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right)$ (13)
$=\frac{1}{4}\left[\text{Var}(\mathbf{W}_{A})+\text{Var}(\mathbf{W}_{B})\right]=\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})$
Mathematical Justification. We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The variances for model A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively. If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$. This shows that, when the variances of the two models are equal, the variance is reduced by a factor of two. More generally, since the average of the variances is always less than or equal to the maximum variance, we have:
$\small\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})\leq\max(\sigma_{A}^{2},\sigma_{B}^{2})$ (14)
Thus, the variance of the averaged model $\overline{\mathbf{W}}$ is always less than or equal to the larger of $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$.
5.3 How Weight Magnitudes and Variance Effects Model Outputs
Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of decision boundary.
According to Wang et al. [32], we have the following property 1 and theorem 1 for the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances:
Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ denotes the set of parameters. Assume the activation function $\phi$ is Lipschitz continuous with constant $L$, and the weight matrices $\mathbf{W}^{(m)}$ are random matrices with independent and identically distributed (i.i.d.) sub-Gaussian entries. Then, for any input $\mathbf{x}$, the output at the $m$-th layer, $\mathbf{y}^{(m)}$, satisfies the following upper bound:
$\small\|\mathbf{y}^{(m)}\|_{2}\leq(L\lambda)^{m}\|\mathbf{x}\|_{2},\text{for }\forall\ \tau>0$ (15)
with a probability of at least $\left(1-2e^{-\tau^{2}}\right)^{m}$, where $(m)$ indexes the neural network layer, with m representing the exponential operation, and $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$ , with $N$ being the maximum number of neurons across all layers, $C_{s}$ representing a constant, and $K_{s}=\max\left\|\mathbf{W}^{(m)}\right\|_{2}$. This property shows that the output norm can grow exponentially with the number of layers $m$, controlled by the Lipschitz constant of the activation function and the bound on the weight matrices.
Theorem 1: Assume the weights in each layer are i.i.d. sub-Gaussian random variables with variance $\sigma_{w}^{2}$, and the biases are with variance $\sigma_{b}^{2}$. Furthermore, let the activation function $\phi$ be Lipschitz continuous with constant $L$. Then, the variance of the output of the final layer, represented by $V(f_{\theta}(x))$, can be bounded by:
$\small\begin{split}V(f_{\theta}(x))\leq&\|x\|_{2}^{2}(L^{2}N)^{m}\prod_{m=1}^{M_{l}}(\sigma^{(m)}_{w})^{2}+L^{2}N\left(\sigma_{b}^{(M_{l})}\right)^{2}+\\ &L^{2}\sum_{m=1}^{M-1}\left\{N\left(\sigma_{b}^{(m)}\right)^{2}\prod_{l=m+1}^{M}\left[L^{2}N\left(\sigma^{(m)}_{w}\right)^{2}\right]\right\},\end{split}$ (16)
where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant. This inequality shows that the output variance grows with the depth $M$ of the network and the variances of the weights and biases.
Property 1 and Theorem 1 show that the increase of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances and vice-versa. As shown in Fig. 5, consider a neural network with input $\mathbf{x}$ and output $\mathbf{y}$, where weights are denoted as $\mathbf{W}$. When model weights have large magnitudes and high variance, predictions become unreliable. This issue can arise from several factors, including task specificity, lack of regularization or improper weight initialization. Together, these factors can drive the weights to unstable values, reducing the model’s ability to generalize effectively. On the other hand, when $\mathbf{W}\to 0$, the output $\mathbf{y}$ will tend to a fixed point determined predominantly by the network’s bias terms, thus becoming largely independent of the input $\mathbf{x}$. Hence, if the magnitude of $\mathbf{W}$ is small, but not identically zero, then the range of outputs for $\mathbf{y}$ will be limited. This reduces the sensitivity of $\mathbf{y}$ to changes in $\mathbf{x}$, effectively shrinking the output space (similar to the L1 or L2 regularization). Such behavior can enhance the robustness of the model, as smaller weight magnitudes reduce the model’s response to minor perturbations in input $\mathbf{x}$. However, this comes with a trade-off: the expressive capacity of the model diminishes as well, potentially leading to poor performance on complex tasks.
{mdframed}
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec. 6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximun magnitudes/variances of individual models before merging. This is because generally the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness. In this case, model-merging is similar to regularization. This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes may improve performance. However, for tasks requiring detailed pattern recognition, lower weight magnitudes may hinder the network’s ability to capture critical pattern distinctions in the data.
6 Experiments
The experimental settings can be found in the supplementary materials.
6.1 The Patterns Contained in Model Weights
(a) Average of all images for each class.
(b) Linear classifiers visualization.
Figure 6: Average of all images for each class and linear classifiers visualization for each class in Tiny ImageNet [13] dataset.
Figure 7: The examples of the first layer weights visualization of Resnet50 with nonlinearity inside the model.
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 clearly shows the patterns of the visual class learned, which are strongly correlated with the average of all images within each class. The template of each class acts as a merged collection of representative images of the respective class. Interestingly, if we merge the weights of first 5 classes in Fig. 1 with the last 5 classes, it gives us the results of Fig. 2. As we analyzed before, the weight-averaged model-merging acts as the Mixup augmentation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that after merging, each element within the combined template becomes noticeably more faded. This phenomenon can also explain the limitation of expressiveness of merged model as described in Sec. 4.
Fig. 6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes. For example, “goldfish” shows a redish center in the averaged images and also shows a redish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visuliazation. However, they are not as clear as for Cifar100. This may due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear. This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [8] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection being similar to the Haar features described in Sec. 3.1.
The visualization for every class in Cifar100 and Tiny ImageNet are put in the supplementary materials.
6.2 Averaging on Weights vs. Averaging on Features
6.2.1 Differences over Different Models
As shown in the Fig. 3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst. We can also discover that greedy soups are generally better than uniform soups due to its mechanism to greedily select the best performing model to merge. It ensures that a good result can be reached by only incorporating the first model. Greedy ensemble on both logits and features also produce results that are comparable to greedy soups.
In general, adding more models into the merging process, leads to better performance for all methods, including model-merging and ensemble. Interestingly, the ViT model, including 3 ViT based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups. This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs). CNNs is more adapted to the learning of local patterns within different scales of images, making it ideal for model-merging. On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model-merging. ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
Figure 8: (a) For Sec. 6.2, performance Gaps between Logits Ensemble and Uniform Soups Across Different Configurations (excluded the ViT models as the scales are too different) on Cifar100. (b) For Sec. 6.3, accuracy vs. Magnification Factor across Methods with 10 models on Cifar100 dataset. (c) For Sec. 6.3, Uniform soup - Logits Ensemble Accuracy vs. Magnification Factor in terms of different model numbers on Cifar100 dataset.
We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (as a function of the number of merged models and different model architectures, but excluding ViT models as their gaps are too large) in Fig. 8. We can see that the more models into the merging process, the larger the gaps between logits ensemble and uniform soups. The gaps are different according to different models, but generally there is down trend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm). This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities. Batch normalization can break this down trend. We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that the greedy soups can alleviate this marginal effect.
6.2.2 Differences over Different Datasets
Table 1: Performance Comparison for VGG19 with 10 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay [31, 16] dataset; while accuracy (0% to 100%) for the rest of datasets.
Datasets Perf Ave Uni Soups Grd Soups Ens Lgt Ens Fts Grd Ens Lgt Grd Ens Fts
Cifar10 92.37 92.90 92.62 93.91 93.81 92.63 92.62
Cifar100 70.34 69.63 70.25 75.36 74.96 70.39 70.32
PathMNIST 90.04 32.42 88.48 92.95 75.17 88.59 88.34
DermaMNIST 74.17 68.18 74.16 75.76 69.73 73.92 73.17
Celeba 92.91 53.19 93.06 93.44 93.47 93.02 93.04
TinyImageNet 59.90 62.31 61.95 63.28 63.53 63.53 63.58
ChestXRay 80.66 79.29 80.43 81.77 81.07 80.43 80.43
For the dataset-wise examination, we perform model merging and ensemble on 7 prevalent datasets with different data sizes. From Tab. 1, we can observe ensemble still generally performs better than merging. Different datasets have different performances and this is due to the characteristics of each dataset. Similarly to the model-wise comparison, with more models in the merging (2-7 models merging results in the supplementary materials), the better the performance for both merging and ensembling. The model ensembling is more stable than model merging in terms of the best and worst performance on different data. For example, on the PathMNIST [37, 38] dataset, uniform soup merging does not work really well compared to the relative performance on the Cifar10/Cifar100 or Tiny ImageNet datasets (that only achieves 40.85% accuracy, which is largely behind model ensemble methods). A similar situation happens on Celeba [15] dataset: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model. These results may be caused by the contradictions between two models’ weights.
6.2.3 Merging Models trained over different Tasks
Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods. Model 1 means VGG19 model trained on Cifar100 dataset; while Model 2 denotes the VGG19 models trained on each different dataset with different task. Ft ens is better than logit ens.
Datasets Size Model1 Model2 Uniform Soups Ens Lgt Ens Features
Cifar10 60,000 70.37 5.84 49.63 69.93 69.93
PathMNIST 107,180 70.37 1.29 24.25 69.88 70.00
DermaMNIST 10,015 70.37 5.77 60.37 69.06 69.36
Celeba 202,599 70.37 1.03 49.33 69.48 69.61
TinyImageNet 100,000 70.37 18.16 61.02 58.40 58.52
ChestXRay 112,120 70.37 15.18 61.38 68.18 68.33
We also explore model merging between different well-trained models on different tasks, as shown in Tab. 2. Generally, merging within Tiny ImageNet (where Model 2 was trained on) has the best performance of 61.02% accuracy, which is better than the ensemble on logits (58.40% accuracy). This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images. Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results does not perform better than model ensemble, as it is too small. On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement. This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data were learned. However, the performance on Celeba is not good. This may be caused by the Celeba dataset only containing human faces, which can be seen as a fine-grained classification task. Due to the uniformity of the data format within Celeba, it contradicts the diverse and rich general computer vision features learned by the model weights on Cifar100.
6.3 Model Predictions and Weight Magnitudes
[…]
Specifically, we visualize the merging of 10 models (merging of 2 to 7 models can be found in the supplementary materials) for model-merging and ensemble towards different magnitude factors, as shown in Fig. 8(b). In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with weight magnification. From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features. But surprisingly, feature ensemble are more robust than logits ensemble across different configurations. Feature ensemble models can even outperform model-merging when weight magnification is not large (e.g., ×$\times 50$). This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances. The “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification. In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model. We can observe the “Ens Features” and “Ens Features*” models perform similarly.
We further visualize the gaps between weight-averaged model-merging and ensemble in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ for Cifar100). The same experiments to visualize the merging of 2 to 10 models for Cifar10 with different magnitude factors, as well as the gaps between weight-averaged model-merging and ensemble, are shown in the supplementary materials. Also, a similar phenomenon is shown on Cifar10, but with different peaks of gaps w.r.t. the magnitudes. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensemble, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g. 4.81% to 6.86% accuracy for 10 models merging). This may be caused the complexity of the Tiny ImageNet task and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization). It will mitigate the weight magnitude changing issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it plays aregularization role. However it does not work forlarger datasets as Tiny ImageNet, because it sacrifices the expressiveness of the models.
Model # Perf Ave xM Perf Uni Soups Grd Soups Ens Lgt Ens Fts Ens Fts* Grd Ens Lgt Grd Ens Fts
2 models 59.87 5.60 6.70 6.99 1.59 6.94 6.45 1.52 6.96
3 models 59.85 5.44 7.15 7.05 0.78 7.38 6.78 0.78 7.21
5 models 59.89 5.07 6.68 6.83 0.30 6.95 6.46 0.35 7.21
7 models 59.90 4.80 6.77 6.80 0.24 6.73 6.26 0.21 6.74
10 models 59.90 4.81 6.86 6.92 0.13 6.92 6.22 0.11 6.95
7 Conclusion
In the paper, we posit weight-averaged model-merging from a template matching perspective by visualizing weight patterns across different datasets. Additionally, our comparison between weight and feature averaging in ensemble strategies revealed distinctions in their impacts across diverse model architectures and datasets, providing valuable insights into when and why each approach can be advantageous. Finally, we found model-merge techniques can provide a prediction that is more stable than model ensemble techniques, particularly under different parameter magnitudes/variances, contributing to increased robustness and generalization. These insights clarify interpretability and critical aspects of model-merging for understanding and applying this technique in various contexts. We hope these findings will encourage further exploration into model-merging methods and inform future development of interpretable and robust model-averaging techniques.
Supplementary Material
8 Experimental Settings
Table 4: The table shows dataset size, class number, and link for each of the dataset. In particular, ChestXRay dataset has 14 labels and each of them is a binary classification task.
Datasets Sizes Class # Links
Cifar10 60,000 10 https://www.cs.toronto.edu/~kriz/cifar.html
Cifar100 60,000 100 https://www.cs.toronto.edu/~kriz/cifar.html
PathMNIST 107,180 9 https://medmnist.com/
DermaMNIST 10,015 7 https://medmnist.com/
CelebA 202,599 2 https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
TinyImageNet 100,000 200 https://huggingface.co/datasets/zh-plus/tiny-imagenet
ChestXRay 112,120 14$\times$2 https://www.kaggle.com/paultimothymooney/chest-xray-pneumonia
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [12] and Tiny ImageNet [13] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size (e.g. 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19), ViT [24, 6] and DeiT [27]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [37, 38], Celeba [15], Tiny ImageNet and ChestXRay [31, 16]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay dataset has 14 labels and each of them is a binary classification task. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay. For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer). Besides these two model ensemble techniques and whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts). In order to exclude the impact from adopting the last layer from any single model, the “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification.
For the experiments of model predictions and weight magnitudes, we conduct experiments on Cifar10, Cifar100 and Tiny ImageNet datasets. To increase the magnitude and the variance of model weights, we magnify the model weights by a constant factor. Formally, when each element in a matrix $A$ is scaled by a constant factor $c$, the variance of the resulting matrix elements increases proportionally to the square of this factor. For example, in the case where $c=100$, the variance of the elements of matrix $A$ becomes $100^{2}\times\text{Var}(A)=10,000\times\text{Var}(A)$.
For all experiments in this section, to keep a fair comparison, we adopt the official splits for each dataset. Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions.
9 More Visualization of Model Weight Patterns
Figure 9: Average of all images within each class (all 100 classes) of Cifar100 dataset.
Figure 10: Linear classifiers visualization for each class (all 100 classes) of Cifar100 dataset.
Similar to the visualization in the main paper, we visualize the average of all images Fig. 9 and linear classifiers for all classes Fig. 10 within each class (all 100 classes) of Cifar100 dataset. We can clearly observe the positive correlation between the average of images in a class and the learned weights.
Figure 11: Merge the linear classifiers for the first 50 classes with the last 50 classes in Cifar100 dataset.
We further visualize the linear classifiers after pair-wise merging between the first 50 classes and the last 50 classes in Cifar100 dataset as Fig. 11. We can see the combinations of two templates, e.g. “motorcycle+woman”.
Figure 12: The first layer weights visualization of Resnet50.
Figure 13: The first layer weights (first 63 conv kernels) visualization of VGG19.
More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 and Fig. 13.
Figure 14: Average of all images within each class (the first 100 classes) of TinyImageNet dataset.
Figure 15: Linear classifiers visualization for each class (the first 100 classes) of TinyImageNet dataset.
Figure 16: Average of all images within each class (the last 100 classes) of TinyImageNet dataset.
Figure 17: Linear classifiers visualization for each class (the last 100 classes) of TinyImageNet dataset.
For Tiny ImageNet, we have Fig. 14 and Fig. 16 for the average of all images within 200 classes in the dataset; also, Fig. 15 and Fig. 17 for the linear classifier visualization of all 200 classes in the dataset. We can observe the correlations between averaged images and the learned patterns.
10 More Experimental Results of Averaging on Weights V.S. Averaging on Features
Figure 18: The 2 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 19: The 3 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 20: The 5 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 21: The 7 model merging/ensemble accuracy comparison between different architectures on Cifar100.
The 2 models, 3 models, 5 models and 7 models for model merging/ensemble accuracy comparison between different architectures on Cifar100 are shown via Fig. 18, Fig. 19, Fig. 20 and Fig. 21, respectively. They all show that the ensemble on logits generally receives the best results and uniform soups receives the worst results. Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
Figure 22: Performance Gaps between Logits Ensemble and Greedy Soups Across Different Configurations (excluded the ViT models as the scales are too different) on Cifar100. It shows the greedy soups can mitigate the marginal effect problem to some extent.
From Fig. 22, we can observe similar with the gaps between ensemble on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be. But differently, the greedy soups break the going down trend with uniform soups.
Table 5: Performance Comparison for VGG19 with 2 Models Across Different Datasets. We adopt averged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
Datasets Perf Ave Uni Soups Grd Soups Ens Lgt Ens Fts Grd Ens Lgt Grd Ens Fts
Cifar10 92.42 92.79 92.68 93.26 93.22 92.68 92.59
Cifar100 70.31 70.06 70.22 72.99 72.76 70.57 70.41
PathMNIST 89.75 40.85 88.41 91.60 82.06 88.61 88.47
DermaMNIST 74.19 71.97 73.57 74.66 72.72 73.87 73.57
Celeba 93.12 89.50 93.05 93.28 93.26 93.08 93.06
TinyImageNet 59.87 60.89 61.40 61.71 61.85 61.72 61.65
ChestXRay 80.52 79.80 80.43 81.14 80.85 80.43 80.43
Table 6: Performance Comparison for VGG19 with 3 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
Cifar10 92.47 92.82 92.65 93.54 93.48 92.75 92.65
Cifar100 70.34 69.80 70.14 73.85 73.49 70.29 70.40
PathMNIST 89.46 45.39 88.65 91.82 80.07 88.41 88.38
DermaMNIST 75.00 69.43 74.01 75.21 72.07 73.47 73.37
Celeba 93.12 77.69 93.12 93.31 93.30 93.01 93.06
TinyImageNet 59.85 61.39 61.55 62.40 62.41 62.60 62.61
ChestXRay 80.56 79.59 80.43 81.31 80.88 80.43 80.43
Table 7: Performance Comparison for VGG19 with 5 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
Cifar10 92.45 92.90 92.65 93.70 93.71 81.95 92.65
Cifar100 70.47 69.73 70.32 74.90 74.18 70.35 70.31
PathMNIST 89.80 51.00 88.58 92.28 79.42 88.47 88.41
DermaMNIST 74.46 69.88 73.32 75.31 72.42 73.57 73.72
Celeba 93.03 60.27 93.06 93.45 93.32 93.06 93.04
TinyImageNet 59.89 61.55 61.91 63.08 63.17 63.29 63.27
ChestXRay 80.57 79.38 80.43 81.46 80.92 80.43 80.43
Table 8: Performance Comparison for VGG19 with 7 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
Cifar10 92.42 92.79 92.65 93.85 93.71 92.76 92.71
Cifar100 70.35 69.79 70.28 74.92 74.82 70.31 70.40
PathMNIST 89.66 38.76 88.44 92.48 79.97 88.50 88.54
DermaMNIST 74.12 69.83 73.67 75.86 70.57 73.82 74.06
Celeba 92.99 49.97 93.07 93.31 93.30 93.05 93.05
TinyImageNet 59.90 62.38 62.12 63.53 63.54 63.66 63.52
ChestXRay 80.60 79.22 80.43 81.61 80.99 80.43 80.43
Also, we show the dataset-wise exploration of 2 models, 3 models, 5 models and 7 models merging/ensembling in Tab. 5, Tab. 6, Tab. 7 and Tab. 8, respectively. The computation consumptions in Floating Point Operations (FLOPs) and Parameter numbers are in Tab. 9.
Table 9: Comparison of FLOPs and Parameters for Various Models. The FLOPs are calculated based on 224 $\times$ 224 for general purpose, e.g. ImageNet. “ViT_Tiny” denotes for “ViT_Tiny_Patch16_224” and “DeiT_Tiny” stands for “DeiT_Tiny_Patch16_224”, respectively.
Models FLOPs (G) Param # (M)
ResNet18 1.82 11.7
ResNet34 3.66 21.8
ResNet50 4.09 25.6
ResNet101 7.83 44.5
ResNet152 11.58 60.2
DenseNet121 2.88 8.0
DenseNet161 7.81 28.7
DenseNet169 3.36 14.3
DenseNet201 4.37 20.0
VGG11 7.63 132.9
VGG11_BN 7.76 132.9
VGG13 11.34 133.0
VGG13_BN 11.47 133.0
VGG16 15.5 138.4
VGG16_BN 15.52 138.4
VGG19 19.6 143.7
VGG19_BN 19.63 143.7
DeiT_Tiny 1.3 5.7
ViT_Tiny 1.3 5.7
ViT_B_16 17.6 86.4
Table 10: Performance Comparison for DeiT on Tiny ImageNet Across Different Configurations. We adopt accuracy (0% to 100%) as the evaluation metric.
Model # Perf Ave Uniform Soups Greedy Soups Ens Logits Ens Features Greedy Ens Logits Greedy Ens Fts
2 models 66.77 61.75 65.99 70.32 69.37 65.99 65.77
3 models 66.32 29.93 65.99 71.56 70.31 65.99 65.77
5 models 66.08 18.71 65.99 72.49 70.86 65.99 65.77
7 models 66.84 13.95 65.99 73.32 72.05 65.99 65.77
10 models 66.79 8.91 65.99 73.93 72.58 65.99 65.77
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models to be merged, the poorer results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features bearly changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
11 More Experimental Results of Model Magnitude Changes
Figure 23: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar100 dataset.
Figure 24: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar100 dataset.
Figure 25: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar100 dataset.
Figure 26: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
The 2 models, 3 models, 5 models and 7 models accuracy vs. magnification factor on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
Figure 27: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset.
Figure 28: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar10 dataset.
Figure 29: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar10 dataset.
Figure 30: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar10 dataset.
Figure 31: Accuracy vs. Magnification Factor across Methods with 10 models on Cifar10 dataset.
Figure 32: Uniform soup - Logits Ensemble Accuracy vs. Magnification Factor across different model numbers on Cifar10 dataset.
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. magnification factor across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons. However, on the Cifar10, the gaps have different peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v2

@@
 In this paper, we address this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
-  • An examination of the patterns contained in the averaged model weights through the visualization of the patterns present in both linear layer as classifiers or inside deep learning models used on multiple datasets.
+  • An examination of the patterns contained in the model weights through the visualization of the patterns present in both linear layers as classifiers or inside deep learning models used on multiple datasets.
 Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
   • An investigation of averaging on weights versus averaging on features strategies allows us to evaluate the relative effectiveness and limitations of model-merging and traditional model ensembles across multiple architectures and datasets, uncovering distinct benefits and behaviors;
-  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of of model regularization.
+  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of model regularization.
 Based on our viewpoints and findings, we conclude with useful proposals and source code that will be publicly available to provide a good reference for the model-merging research community.
@@
 Yu et al.
-[39] combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
-Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited.
+[39] combine models trained on domain-specific data to create a domain-agnostic model merging that generalizes well across diverse distributions.
+Despite these advances, research into the underlying mechanisms that make model merging effective remains limited.
 This gap in understanding serves as the motivation for our paper.
@@
 In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching.
-Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with inner product which is a widely explored operation in neural networks.
+Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks.
 By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
@@
 During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
-In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in the Fig.
+In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in Fig.
 6.1.
@@
 The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space.
-This also explains why recent research [5, 41, 10] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
-This concept extends to model-merging, where multiple models or variations of a model are combined.
+This also explains why recent research [5, 41, 10, 44] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
+This concept extends to model merging, where multiple models or variations of a model are combined.
 In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model.
-The model after merging can work more effectively than individuals because they represent a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
+The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
 3.2 Formal Representation from the Perspective of Template Matching
@@
 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights.
-However, the averaging on weights can show relatively good performance without introducing extra inference computational cost.
-Hence, this forms a trade-off between computation usage and model performance when deciding between model-merging or model ensemble.
+However, the averaging on weights can show relatively good performance without introducing extra inference computational costs.
+Hence, this forms a trade-off between computation usage and model performance when deciding between model merging or model ensemble.
 5 The Model Predictions and Weight Magnitudes
-Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes on Cifar100.
+Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes with VGG19 model on Cifar100.
 The values on the X and Y axis represent the accuracies.
@@
 Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$.
-This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound equal to the maximum magnitude of the original weights, but in most cases it will be less than that maximum.
+This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound is equal to the maximum magnitude of the original weights, but in most cases, it will be less than that maximum.
 Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
@@
 We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$.
-The variances for model A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively.
+The variances for models A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively.
 If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$.
@@
 5.3 How Weight Magnitudes and Variance Effects Model Outputs
-Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of decision boundary.
+Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of the decision boundary.
 According to Wang et al. [32], we have the following property 1 and theorem 1 for the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances:
@@
 [backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec.
-6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximun magnitudes/variances of individual models before merging.
-This is because generally the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness.
-In this case, model-merging is similar to regularization.
+6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging.
+This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness.
+In this case, model merging is similar to regularization.
 This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements.
@@
 The template of each class acts as a merged collection of representative images of the respective class.
-Interestingly, if we merge the weights of first 5 classes in Fig.
+Interestingly, if we merge the weights of the first 5 classes in Fig.
 1 with the last 5 classes, it gives us the results of Fig.
@@
 For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle.
-We can also observe that after merging, each element within the combined template becomes noticeably more faded.
-This phenomenon can also explain the limitation of expressiveness of merged model as described in Sec.
+We can also observe that after merging, each element within the combined template noticeably fades.
+This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec.
 4.
@@
 6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes.
-For example, “goldfish” shows a redish center in the averaged images and also shows a redish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visuliazation.
+For example, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visualization.
 However, they are not as clear as for Cifar100.
-This may due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear.
+This may be due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear.
 This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100.
@@
 7 (VGG model weight visualization can be found in the supplementary materials).
-We can see the patterns of texture and edges for the basic level geometric feature detection being similar to the Haar features described in Sec.
+We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec.
 3.1.
-The visualization for every class in Cifar100 and Tiny ImageNet are put in the supplementary materials.
-6.2 Averaging on Weights vs. Averaging on Features
-6.2.1 Differences over Different Models
-As shown in the Fig.
+The visualization for every class in Cifar100 and Tiny ImageNet is put in the supplementary materials.
+6.2 Weights vs. Features Averaging
+6.2.1 Observations over Different Models
+As shown in Fig.
 3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst.
-We can also discover that greedy soups are generally better than uniform soups due to its mechanism to greedily select the best performing model to merge.
+We can also discover that greedy soups are generally better than uniform soups due to their mechanism to greedily select the best performing model to merge.
 It ensures that a good result can be reached by only incorporating the first model.
-Greedy ensemble on both logits and features also produce results that are comparable to greedy soups.
-In general, adding more models into the merging process, leads to better performance for all methods, including model-merging and ensemble.
-Interestingly, the ViT model, including 3 ViT based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups.
+The greedy ensemble on both logits and features also produces results that are comparable to greedy soups.
+In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensemble.
+Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups.
 This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs).
-CNNs is more adapted to the learning of local patterns within different scales of images, making it ideal for model-merging.
-On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model-merging.
+CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for model merging.
+On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model merging.
 ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
@@
 8.
-We can see that the more models into the merging process, the larger the gaps between logits ensemble and uniform soups.
-The gaps are different according to different models, but generally there is down trend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm).
+We can see that the more models in the merging process, the larger the gaps between logits ensemble and uniform soups.
+The gaps are different according to different models, but generally, there is a downtrend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm).
 This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities.
 Batch normalization can break this downtrend.
-We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that the greedy soups can alleviate this marginal effect.
-6.2.2 Differences over Different Datasets
+We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that greedy soups can alleviate this marginal effect.
+6.2.2 Observations over Different Datasets
 Table 1: Performance Comparison for VGG19 with 10 Models Across Different Datasets.
@@
 Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model.
-These results may be caused by the contradictions between two models’ weights.
+These results may be caused by the contradictions between the two models’ weights.
 6.2.3 Merging Models trained over different Tasks
 Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods.
-Model 1 means VGG19 model trained on Cifar100 dataset; while Model 2 denotes the VGG19 models trained on each different dataset with different task.
-Ft ens is better than logit ens.
+Model 1 is the VGG19 model trained on the Cifar100 dataset, while Model 2 denotes the VGG19 model trained on each different dataset with different tasks.
+Feature ensemble is better than logits ensemble.
 Datasets Size Model1 Model2 Uniform Soups Ens Lgt Ens Features
@@
 This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images.
-Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results does not perform better than model ensemble, as it is too small.
+Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results do not perform better than the model ensemble, as it is too small.
 On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement.
-This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data were learned.
+This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data was learned.
 However, the performance on Celeba is not good.
@@
 From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features.
-But surprisingly, feature ensemble are more robust than logits ensemble across different configurations.
-Feature ensemble models can even outperform model-merging when weight magnification is not large (e.g., ×$\times 50$).
+But surprisingly, feature ensembles are more robust than logits ensembles across different configurations.
+Feature ensemble models can even outperform model merging when weight magnification is not large (e.g., ×$\times 50$).
 This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances.
-The “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification.
+The “Ens Features” model uses the merge of the last fully connected (FC) layers from all merged models for classification.
 In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model.
@@
 3.
-From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent.
-It can also outperform logit ensemble, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g.
+From the table, we can observe that model merging can still mitigate the weight magnifying problem to a certain extent.
+It can also outperform logit ensembles, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g.
 4.81% to 6.86% accuracy for 10 models merging).
-This may be caused the complexity of the Tiny ImageNet task and as we discussed in the Sec.
+This may be caused by the complexity of the Tiny ImageNet task and as we discussed in the Sec.
 5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization).
@@
 This approach works for smaller data (Cifar10/Cifar100) as it plays a regularization role.
-However it does not work forlarger datasets as Tiny ImageNet, because it sacrifices the expressiveness of the models.
+However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
 Model # Perf Ave xM Perf Uni Soups Grd Soups Ens Lgt Ens Fts Ens Fts* Grd Ens Lgt Grd Ens Fts
@@
 7 Conclusion
-In the paper, we posit weight-averaged model-merging from a template matching perspective by visualizing weight patterns across different datasets.
-Additionally, our comparison between weight and feature averaging in ensemble strategies revealed distinctions in their impacts across diverse model architectures and datasets, providing valuable insights into when and why each approach can be advantageous.
-Finally, we found model-merge techniques can provide a prediction that is more stable than model ensemble techniques, particularly under different parameter magnitudes/variances, contributing to increased robustness and generalization.
-These insights clarify interpretability and critical aspects of model-merging for understanding and applying this technique in various contexts.
-We hope these findings will encourage further exploration into model-merging methods and inform future development of interpretable and robust model-averaging techniques.
+In this paper, we thoroughly examined the weight-averaged model-merging approach through the lens of template matching by visualizing weight patterns across various datasets.
+Our comparison of weight and feature averaging within ensemble and merging methods revealed notable differences in their effects on different model architectures and datasets, offering insights into when and why each approach is beneficial.
+Furthermore, we demonstrate that model merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization.
+These findings clarify the interpretability and key aspects of model-merging, providing a deeper understanding of its application across diverse contexts.
+We anticipate that these insights will stimulate further research into model merging strategies and support the development of more interpretable and robust model-averaging methodologies.
 Supplementary Material
@@
 The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1.
-While visualization, the weights are wrapped back to the image size (e.g.
-100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
+While visualization, the weights are wrapped back to the image size.
 In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging.
 We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances.
-For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19), ViT [24, 6] and DeiT [27].
+For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19, including with and without batch normalization), ViT [24, 6] and DeiT [27].
 For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [37, 38], Celeba [15], Tiny ImageNet and ChestXRay [31, 16].
@@
 For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer).
-Besides these two model ensemble techniques and whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts).
-In order to exclude the impact from adopting the last layer from any single model, the “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification.
+Besides these two model ensemble techniques, whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts).
+In order to exclude the impact of adopting the last layer from any single model, the “Ens Features” model uses the merging of the last fully connected (FC) layers from all merged models for classification.
 For the experiments of model predictions and weight magnitudes, we conduct experiments on Cifar10, Cifar100 and Tiny ImageNet datasets.
@@
 For example, in the case where $c=100$, the variance of the elements of matrix $A$ becomes $100^{2}\times\text{Var}(A)=10,000\times\text{Var}(A)$.
-For all experiments in this section, to keep a fair comparison, we adopt the official splits for each dataset.
-Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions.
+For all experiments in the paper, to keep fair comparisons, we adopt the official splits for each dataset.
+Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions (learning rates, step strategies and optimizers, etc).
 9 More Visualization of Model Weight Patterns
-Figure 9: Average of all images within each class (all 100 classes) of Cifar100 dataset.
-Figure 10: Linear classifiers visualization for each class (all 100 classes) of Cifar100 dataset.
+Figure 9: Average of all images within each class of Cifar100 dataset (all 100 classes).
+Figure 10: Linear classifiers visualization for each class of Cifar100 dataset (all 100 classes).
 Similar to the visualization in the main paper, we visualize the average of all images Fig.
@@
 10 within each class (all 100 classes) of Cifar100 dataset.
-We can clearly observe the positive correlation between the average of images in a class and the learned weights.
+We can easily and clearly observe the strong correlation between the average of images in a class and the learned weights.
 Figure 11: Merge the linear classifiers for the first 50 classes with the last 50 classes in Cifar100 dataset.
@@
 Figure 13: The first layer weights (first 63 conv kernels) visualization of VGG19.
-More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 and Fig. 13.
-Figure 14: Average of all images within each class (the first 100 classes) of TinyImageNet dataset.
-Figure 15: Linear classifiers visualization for each class (the first 100 classes) of TinyImageNet dataset.
-Figure 16: Average of all images within each class (the last 100 classes) of TinyImageNet dataset.
-Figure 17: Linear classifiers visualization for each class (the last 100 classes) of TinyImageNet dataset.
+More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g.
+Fig.
+12 for ResNet50 and Fig.
+13 for VGG19.
+For ResNet50 kernel visualization, we can observe differences or similarities between each learned patterns.
+For instance, Kernel 5 and kernel 9 are similar, which are responsible for detecting similar features and thus yield a higher measurement (e.g.
+inner product) between the two kernels; while Kernel 9 is very different with Kernel 45.The patterns inside ResNet50 is easier to be observed than VGG19, as ResNet50 has larger kernel size for the 1st layer.
+Figure 14: Average of all images within each class of TinyImageNet dataset (the first 100 classes).
+Figure 15: Linear classifiers visualization for each class of TinyImageNet dataset (the first 100 classes).
+Figure 16: Average of all images within each class of TinyImageNet dataset (the last 100 classes).
+Figure 17: Linear classifiers visualization for each class of TinyImageNet dataset (the last 100 classes).
 For Tiny ImageNet, we have Fig.
@@
 10 More Experimental Results of Averaging on Weights V.S. Averaging on Features
-Figure 18: The 2 model merging/ensemble accuracy comparison between different architectures on Cifar100.
-Figure 19: The 3 model merging/ensemble accuracy comparison between different architectures on Cifar100.
-Figure 20: The 5 model merging/ensemble accuracy comparison between different architectures on Cifar100.
-Figure 21: The 7 model merging/ensemble accuracy comparison between different architectures on Cifar100.
-The 2 models, 3 models, 5 models and 7 models for model merging/ensemble accuracy comparison between different architectures on Cifar100 are shown via Fig.
+Figure 18: The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+Figure 19: The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+Figure 20: The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+Figure 21: The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+The 2 models, 3 models, 5 models and 7 models for model merging/ensembling accuracy comparison between different architectures on Cifar100 are shown via Fig.
 18, Fig.
@@
 21, respectively.
-They all show that the ensemble on logits generally receives the best results and uniform soups receives the worst results.
+They all show that the models with ensembling on logits generally receive the best results and uniform soups receive the worst results.
 Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
 Figure 22: Performance Gaps between Logits Ensemble and Greedy Soups Across Different Configurations (excluded the ViT models as the scales are too different) on Cifar100.
-It shows the greedy soups can mitigate the marginal effect problem to some extent.
-From Fig.
-22, we can observe similar with the gaps between ensemble on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be.
-But differently, the greedy soups break the going down trend with uniform soups.
+Fig.
+22 shows the performance gaps between ensembling on logits with greedy soups.
+From it, we can observe similar phenomenon with the gaps between ensembling on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be.
+But differently, the greedy soups break the going down trend with uniform soups as described in the main paper.
 Table 5: Performance Comparison for VGG19 with 2 Models Across Different Datasets.
-We adopt averged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
+For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
 Datasets Perf Ave Uni Soups Grd Soups Ens Lgt Ens Fts Grd Ens Lgt Grd Ens Fts
@@
 Table 6: Performance Comparison for VGG19 with 3 Models Across Different Datasets.
-We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
+For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
 Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
@@
 Table 7: Performance Comparison for VGG19 with 5 Models Across Different Datasets.
-We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
+For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
 Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
@@
 Table 8: Performance Comparison for VGG19 with 7 Models Across Different Datasets.
-We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
+For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
 Datasets Perf Ave Uni Soups Grd Soups Ens Logits Ens Fts Grd Ens Logits Grd Ens Fts
@@
 9.
-Table 9: Comparison of FLOPs and Parameters for Various Models.
+Table 9: Comparison of FLOPs and Parameter Numbers for Various Models.
 The FLOPs are calculated based on 224 $\times$ 224 for general purpose, e.g.
@@
 We can also see the more models to be merged, the poorer results from uniform soups can get.
-Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features bearly changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
+Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
 11 More Experimental Results of Model Magnitude Changes
@@
 We can observe similar phenomenons.
-However, on the Cifar10, the gaps have different peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).
+However, on the Cifar10, the gaps have different a peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v3

@@
 Abstract
-Weight-averaged model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without fine-tuning or retraining.
+Model-merging has emerged as a powerful approach in deep learning, capable of enhancing model performance without any training.
 However, the underlying mechanisms that explain its effectiveness remain largely unexplored.
-In this paper, we investigate this technique from three novel perspectives to provide deeper insights into how and why weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns; (2) we investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales.
+In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns and that is the essential why model-merging can work; (2) we mathematically and empirically investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales.
 Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process.
+The code is available at https://github.com/billhhh/Rethink-Merge.
 1 Introduction
-Model-merging can combine multiple independently trained models into a single, efficient model without introducing extra inference computations.
-This approach has been applied in contexts such as natural language processing [36], computer vision [14], and their sub-areas (e.g.
-federated learning [30], distillation [11], adversarial learning [42, 3] and large language model [20, 43]), where the ability to merge models effectively can lead to improved model performance, efficiency, and flexibility in deployment.
-Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective, especially weight-averaged model-merging [33] (e.g.
-Uniform Soups and Greedy Soups) as the mainstream techniques.
-Without a clear understanding of the underlying mechanisms of model-merging, researchers and practitioners may lack guidance on when and how to apply it in a principled way.
-This research gap motivates the need to systematically explore the underlying patterns and behaviors in weight-averaged model-merging.
-In this paper, we address this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
+Model-merging can combine multiple independently trained models into a single, effective model without introducing extra inference computations.
+This family of approaches has been applied in contexts such as natural language processing [38], computer vision [16], and their sub-areas (e.g.
+federated learning [32], distillation [13], adversarial learning [44, 2] and large language model [22, 45]), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment.
+Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective.
+This research gap motivates the need to innovatively explore the underlying principles of why and how the family of weight-averaged model-merging methods can work.
+In this paper, we bridge this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
   • An examination of the patterns contained in the model weights through the visualization of the patterns present in both linear layers as classifiers or inside deep learning models used on multiple datasets.
@@
   • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of model regularization.
-Based on our viewpoints and findings, we conclude with useful proposals and source code that will be publicly available to provide a good reference for the model-merging research community.
+Based on our viewpoints and findings, we conclude with useful proposals to provide a good reference for the community.
 2 Related Work
@@
 (b) Linear classifiers visualization.
-Figure 1: Average of all images for each class and linear classifiers visualization for each class in Cifar100 [12] dataset. The detailed experimental analysis is in Sec. 6.1.
+Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 [14] dataset (b).
+We train a single linear classifier for all classes then wrap it back to the size of input images.
+The detailed experimental analysis is in Sec.
+6.1.
 Figure 2: Merge the linear classifiers for the first 5 classes (within the above 10 classes) with the last 5 classes on Cifar100.
 It interprets the reason why the model-merging techniques can work as they essentially operate the templates contained in model weights.
-In recent years, model-merging [26, 17, 21, 4, 2, 19, 23, 35] has attracted growing interest.
-This approach has led to a range of methods focused on averaging or merging weights from independently trained models to enhance model performance or generalization.
+In recent years, model-merging [28, 23, 3, 21, 25, 37] has attracted growing interest.
+This approach has led to a range of methods focused on merging weights from independently trained models to enhance model performance or generalization.
 For example, Wortsman et al.
-[33] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup”.
+[35] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference.
 Further, Yadav et al.
-proposed Ties-Merging [34] based on parameter averaging by considering conflicts between the model weights during training, allowing for smoother parameter alignment across models for merging.
-Some other approaches [1, 25] address the problem from a different perspective, focusing on finding weight merging operations via the similarity analyses of model features.
+proposed Ties-Merging [36] based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models.
+Some other approaches [1, 6, 12] address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins.
+Zipit [27] takes one step ahead from Git Re-Basin [1] to align weights from inter-model and intra-model based on the alignment of features between merging models.
 Yu et al.
-[39] combine models trained on domain-specific data to create a domain-agnostic model merging that generalizes well across diverse distributions.
-Despite these advances, research into the underlying mechanisms that make model merging effective remains limited.
+[41] proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
+This approach ensures that domain knowledge is preserved while enhancing the model’s ability to generalize across different tasks.
+Beyond model-merging, other techniques, such as mode connectivity path ensembles [8], analyze pathways in the parameter space between models.
+By identifying optimal modes along these paths, these models help improve the performance after merging.
+Fisher Merging [19] selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
+Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited.
 This gap in understanding serves as the motivation for our paper.
-To focus on a specific scope, we aim to offer a more thorough analysis of weight-averaged model-merging than the study in [33], which serves as the foundation work for the area.
+To focus on a specific scope, we aim to offer a thorough analysis from mathematical and empirical perspectives of weight-averaged model-merging different from the study in [35], which serves as the foundation work for the area.
 3 The Patterns Contained in Model Weights
-In this section, we will explain the effectiveness of model soups, built by averaging model weights, via the analysis of their weights from the perspective of template matching.
-Specifically, we explain why the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks.
-By treating the weights as templates, we can understand that operations in weight space as efforts to manipulate feature representations.
+In this section, we will explain the reason why does the weight-averaged model-merging work via the analysis of their weights from the perspective of template matching (an example is shown Fig.
+1 and Fig.
+2 for a linear model; for nonlinear models, visualizations can be found in the supplementary materials).
+Specifically, we explain the reason that the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks.
+By treating the weights as templates, it is understandable that operations in weight space as efforts to manipulate feature representations.
 3.1 Connections Between Inner Product and Template Matching
-The weights of the model can be interpreted as templates or prototypes for the respective class activation.
-A couple of decades ago, Adaboost [7] adopted heuristically designed patterns such as HAAR features [29], which are simple, hand-crafted filters that capture edges, lines, and other low-level image characteristics through weighted sums and differences in pixel intensities.
+The weights of the model can be interpreted as templates or prototypes for the respective class activations and it has a deep connection with machine learning.
+A couple of decades ago, Adaboost [7] contained heuristically designed patterns such as HAAR features [31], which are simple hand-crafted filters that capture edges / lines / other low-level image features through weighted sums and differences in pixel intensities.
 During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
-In a linear classifier, the learned weight matrix serves as simplified templates representing each class within the data, as shown in Fig.
+In a linear classification, once the weights are learned, we wrap it back to have the same size with the input images.
+Take Cifar100 for instance, each image within the dataset is 32$\times$32$\times$3, if we flatten it and it will become a 3072$\times$1 vector; the weight matrix size is 100$\times$3072 without considering bias; in this case, each row of the weight matrix corresponds to each class to be classified and thus each row vector (1$\times$3072) can be wrapped back to 32$\times$32$\times$3 with the same size of the original images.
+The learned weight matrix serves as simplified templates that represent each class within the data, as shown in Fig.
 6.1.
 A key mechanism enabling these templates to operate effectively is the use of inner products to apply them to the input data.
-When an image is presented to the model, its activation is determined by the inner product between the image’s feature representation and each class’s template.
-The class whose template yields the highest activation is selected as the predicted class, indicating that the template matches the input image most closely.
-This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different “important” features that contribute the most towards the objectives.
-This inner product computation is essentially a similarity measurement.
-In this context, the averaging operation in the weight space represents a linear combination of weight matrices (shown as Fig.
-2), which can be seen as Mixup [40] operation in the weight space.
-The linear combinations of samples in the weight space has a similar effect as in the feature space, e.g., Word2vec [18] as the arithmetic can be applied on the embedding space.
-This also explains why recent research [5, 41, 10, 44] leverages “model arithmetic” of multiple models in the weight space to merge models effectively.
-This concept extends to model merging, where multiple models or variations of a model are combined.
-In such scenarios, the combined model retains the diverse templates or patterns learned from each ingredient model.
-The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input, the merged model will provide a good match.
+When an image is presented to the model, its activation is determined by the inner product between the image features and per class weight vector.
+For each class, which weight vector yields the highest activation is selected as the predicted class of that vector, indicating that the “template” aligns (or similar to) the input image most closely.
+This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g.
+in the lower layer, low level features, e.g.
+edge / line features, are tend to be learned).
+The averaging operation in the weight space represents as Fig.
+2, which can be seen as the Mixup [42] operation in the weight space.
+Therefore, the linear combinations in the weight space has a similar effect as in the feature space.
+This also explains why recent research [4, 43, 11, 46] applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec [20].
+Extending to model-merging, the combined model retains the diverse patterns learned from each ingredient model.
+The model after merging can work more effectively than individuals because it represents a broader set of learned templates, increasing the chances that for any given input the merged model provides a good match.
+Also, it may provide an effect of regularization by merging different sophistically learned patterns, which enhances the model robustness and generalization ability.
 3.2 Formal Representation from the Perspective of Template Matching
-For a linear classifier, each row of the weight vector $\mathbf{w}\in\mathbb{R}^{n}$ can be interpreted as a template for a particular class, and the input vector is denoted as $\mathbf{x}\in\mathbb{R}^{n}$.
+As the inner product computation adopted in linear classification is essentially a similarity measurement, so for a linear classifier, each row of the weight vector $\mathbf{w}\in\mathbb{R}^{n}$ can be interpreted as a template for a particular class.
+The input vector is denoted as $\mathbf{x}\in\mathbb{R}^{n}$.
 The inner product between $\mathbf{w}$ and $\mathbf{x}$ measures the alignment between them:
@@
 On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned.
-This mechanism also applies for the linear layers within deep learning models.
+This mechanism also applies for the layers within deep learning models.
+The visualization of deep models can be found in the supplementary materials.
 {mdframed}
-[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 1: From the analysis above and the experimental study shown in Sec.
-6.1, since weight matrices contain generalizable template information, if the merged model’s performance is suboptimal, more focus should be directed toward resolving conflicts within the individual learned patterns in the weights.
+[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 1: The analyses above and the experimental study in Sec.
+6.1 show weight matrices contain generalizable template information.
+If in practice, the merged model’s performance is suboptimal, more attention should be directed toward resolving conflicts within the individual learned patterns in the weights.
 4 Averaging on Weights vs. Averaging on Features
@@
 The resulting output feature vector is then computed as:
-$\small\overline{\mathbf{h}}=\overline{\mathbf{W}}\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}$ (3)
+$\small\overline{\mathbf{h}}=\overline{\mathbf{W}}\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}=\frac{1}{2}\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x}\right).$ (3)
 Since both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ represent linear transformations, the overall transformation remains linear.
@@
 $\small\overline{\mathbf{h}}^{\prime}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x}\right).$ (5)
-In the context of linear operations, the methods in (3) and (5) are equivalent and yield the same result $\overline{\mathbf{h}}=\overline{\mathbf{h}}^{\prime}$
-4.2 Impact of Non-linear Activation Functions
-The equivalence between averaging on weights or on features breaks when non-linear activation functions $\phi(\cdot)$, such as ReLU or Sigmoid, are introduced.
+In the context of linear operations, the methods in (3) and (5) are equivalent and yield the same result $\overline{\mathbf{h}}=\overline{\mathbf{h}}^{\prime}$.
+4.2 Non-linear Model Scenarios
+The equivalence between averaging on weights or on features breaks when non-linear activation functions $\phi(\cdot)$, such as ReLU or Sigmoid, are involved.
 In the case of averaging on weights, the non-linear activation is applied after averaging the weight matrices:
-$\small\overline{\mathbf{h}}=\phi\left(\overline{\mathbf{W}}\mathbf{x}\right)=\phi\left[\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}\right].$ (6)
-Here, the non-linearity is applied to the weighted average of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$.
+$\small\overline{\mathbf{h}}=\phi\left(\overline{\mathbf{W}}\mathbf{x}\right)=\phi\left[\frac{1}{2}\left(\mathbf{W}_{1}+\mathbf{W}_{2}\right)\mathbf{x}\right]=\phi\left[\frac{1}{2}(\mathbf{W}_{1}\mathbf{x}+\mathbf{W}_{2}\mathbf{x})\right].$ (6)
 When averaging on features, the activation function is applied to each output feature independently:
@@
 The final output is then the average of these activated features:
-$\small\overline{\mathbf{h}}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left[\phi\left(\mathbf{W}_{1}\mathbf{x}\right)+\phi\left(\mathbf{W}_{2}\mathbf{x}\right)\right].$ (8)
-Hence, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models will no longer hold for non-linear models.
-If we take ReLU for instance, each weight matrix retains its positive contribution through non-linear activation functions, while squashing to zero the negative contributions.
-As a result, averaging on features tends to preserve more information than averaging on weights, especially in scenarios where the transformations $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit diverse activations, such as when $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ have different signs, where their average may tend towards zero, leading to potential loss of information after the non-linear activation.
-The key difference between the two approaches lies in how they interact with non-linear activation functions.
+$\small\overline{\mathbf{h}}^{\prime}=\frac{1}{2}\left(\mathbf{h}_{1}+\mathbf{h}_{2}\right)=\frac{1}{2}\left[\phi\left(\mathbf{W}_{1}\mathbf{x}\right)+\phi\left(\mathbf{W}_{2}\mathbf{x}\right)\right].$ (8)
+In such a case, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models may not hold for non-linear models.
+Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to large loss of information after the non-linear activation.
 Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect.
-However, this smoothing may also lead to a loss of fine-grained feature details, especially after non-linear activation, as illustrated in Eq.
-6.
-In contrast, averaging on features allows each weight matrix to independently influence the final output after activation, also better adapting to individual data samples, preserving more variability and potentially enhancing the model’s expressive capacity.
-However, model ensemble will increase the computational cost linearly with the number of ensembled models; while model-merging will not.
+However, this smoothing may also lead to a loss while extracting fine-grained details.
+In contrast, averaging on features allows each weight matrix to independently influence the final output after activation.
+However, averaging on features (model ensemble) increases the computational costs linearly with the number of ensembled models, while averaging on weights (model-merging) does not suffer.
 The FLOPs of models considered in the paper can be found in the supplementary materials.
 {mdframed}
-[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analysis above and the experimental study shown in Sec.
+[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 2: From the analyses above and the experimental study shown in Sec.
 6.2, averaging on features can generally produce stronger expressive ability than averaging on weights.
-However, the averaging on weights can show relatively good performance without introducing extra inference computational costs.
-Hence, this forms a trade-off between computation usage and model performance when deciding between model merging or model ensemble.
+However, the averaging on weights can show promising performance without introducing extra inference computational overhead.
+Hence, this forms a trade-off between computation cost and model performance while deciding between model-merging or model ensembling.
 5 The Model Predictions and Weight Magnitudes
-Figure 4: Scatter plot comparing the accuracies of uniform soup and logits ensemble over different weights magnitudes with VGG19 model on Cifar100.
+Figure 4: Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100.
 The values on the X and Y axis represent the accuracies.
@@
 6.3.
-In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors.
-This helps us understand how the model-merging works as a similar way of regularization.
+In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how robust model-merging works.
 5.1 Analysis of Weight Magnitude After Averaging
-In this section, we analyze the effect of averaging the weights of two independent models, denoted as $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$, on their respective magnitudes.
-We define the maximum magnitude of the weights from both models as $M=\max(|\mathbf{W}_{A}|,|\mathbf{W}_{B}|)$.
+We start from analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes.
+We define the maximum entry of both matrices as $M=\max(\mathbf{W}_{1},\mathbf{W}_{2})=\max_{i,j}(|(\mathbf{W}_{1})_{i,j}|,|(\mathbf{W}_{2})_{i,j}|)$.
 When we compute the averaged weights, we obtain:
-$\small\overline{\mathbf{W}}=\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}.$ (9)
-We seek to understand the relationship between the magnitude of the averaged weights $|\overline{\mathbf{W}}|$ and the maximum magnitude $M$.
+$\small\overline{\mathbf{W}}=\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}.$ (9)
+We seek to understand the relationship between the max entry of the averaged weights $\overline{\mathbf{W}}$ and $M$.
 Mathematical Justification. Using the triangle inequality [30], we establish:
-$\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right\|\leq\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2},$ (10)
+$\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right\|\leq\frac{\|\mathbf{W}_{1}\|+\|\mathbf{W}_{2}\|}{2},$ (10)
 […]
-$\small\frac{\|\mathbf{W}_{A}\|+\|\mathbf{W}_{B}\|}{2}\leq\max(\|\mathbf{W}_{A}\|,\|\mathbf{W}_{B}\|)=M.$ (11)
-Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$.
-This analysis confirms that averaging the weights of two models results in a new weight whose magnitude’s upper bound is equal to the maximum magnitude of the original weights, but in most cases, it will be less than that maximum.
-Our analysis demonstrates that the operation of averaging the weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ typically results in a reduction of the maximum weight magnitude.
-5.2 Variance Reduction in Model Weight Averaging
-Consider two models, each with corresponding weight vectors $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$.
-Let us denote the means and variances of the weight distributions for $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ as $\mathbb{E}[\mathbf{W}_{A}]=\mu_{A}$, $\text{Var}(\mathbf{W}_{A})=\sigma_{A}^{2}$ and $\mathbb{E}[\mathbf{W}_{B}]=\mu_{B}$, $\text{Var}(\mathbf{W}_{B})=\sigma_{B}^{2}$.
-The average of the weights of models A and B is defined as $\overline{\mathbf{W}}=(\mathbf{W}_{A}+\mathbf{W}_{B})/2$
-Our goal is to compute the variance of $\overline{\mathbf{W}}$ and compare it to the variances of $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$.
+$\small\frac{\|\mathbf{W}_{1}\|+\|\mathbf{W}_{2}\|}{2}\leq\max(\mathbf{W}_{1},\mathbf{W}_{2})=M.$ (11)
+Thus, it follows that $\|\overline{\mathbf{W}}\|\leq M$ combining with Eq.
+10.
+It confirms that averaging the weights results in a new weight whose entry’s upper bound is equal to the maximum entry of the original weights.
+Thus, the operation of averaging the weights $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ typically results in a reduction of the maximum weight magnitude.
+This conclusion also extends to complex deep neural networks.
+5.2 Weight Variance Reduces while Averaging
+We denote the means and variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ as $\mathbb{E}(\mathbf{W}_{1})=\mu_{1}$, $\text{Var}(\mathbf{W}_{1})=\sigma_{1}^{2}$ and $\mathbb{E}(\mathbf{W}_{2})=\mu_{2}$, $\text{Var}(\mathbf{W}_{2})=\sigma_{2}^{2}$.
+Our goal is to compute the variance of $\overline{\mathbf{W}}$ and compare it to the variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$.
 The expectation of the averaged weights $\overline{\mathbf{W}}$ is given by:
-$\small\mathbb{E}[\overline{\mathbf{W}}]=\mathbb{E}\left[\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right]=\frac{\mathbb{E}[\mathbf{W}_{A}]+\mathbb{E}[\mathbf{W}_{B}]}{2}=\frac{\mu_{A}+\mu_{B}}{2}$ (12)
+$\small\mathbb{E}(\overline{\mathbf{W}})=\mathbb{E}\left(\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right)=\frac{\mathbb{E}(\mathbf{W}_{1})+\mathbb{E}(\mathbf{W}_{2})}{2}=\frac{\mu_{1}+\mu_{2}}{2}$ (12)
 For the variance of $\overline{\mathbf{W}}$, we have:
-$\small\text{Var}(\overline{\mathbf{W}})=\text{Var}\left(\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right)$ (13)
-$=\frac{1}{4}\left[\text{Var}(\mathbf{W}_{A})+\text{Var}(\mathbf{W}_{B})\right]=\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})$
+$\small\text{Var}(\overline{\mathbf{W}})=\text{Var}\left(\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right)$ (13)
+$=\frac{1}{4}\left[\text{Var}(\mathbf{W}_{1})+\text{Var}(\mathbf{W}_{2})\right]=\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})$
 Mathematical Justification.
-We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$.
-The variances for models A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively.
-If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$.
-This shows that, when the variances of the two models are equal, the variance is reduced by a factor of two.
-More generally, since the average of the variances is always less than or equal to the maximum variance, we have:
-$\small\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})\leq\max(\sigma_{A}^{2},\sigma_{B}^{2})$ (14)
-Thus, the variance of the averaged model $\overline{\mathbf{W}}$ is always less than or equal to the larger of $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$.
+We now compare the variance of $\overline{\mathbf{W}}$ to the variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$.
+If $\sigma_{1}=\sigma_{2}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$.
+This shows that, when the variances of the two models are the same, the variance is reduced by a factor of two.
+More generally, we have:
+$\small\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\max(\sigma_{1}^{2},\sigma_{2}^{2})$ (14)
+Thus, the variance of the averaged model $\overline{\mathbf{W}}$ is always less than or equal to the larger member between $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$.
 5.3 How Weight Magnitudes and Variance Effects Model Outputs
 Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of the decision boundary.
-According to Wang et al. [32], we have the following property 1 and theorem 1 for the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances:
-Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ denotes the set of parameters.
+We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  [34]:
+Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ is the parameters.
 Assume the activation function $\phi$ is Lipschitz continuous with constant $L$, and the weight matrices $\mathbf{W}^{(m)}$ are random matrices with independent and identically distributed (i.i.d.) sub-Gaussian entries.
@@
 with a probability of at least $\left(1-2e^{-\tau^{2}}\right)^{m}$, where $(m)$ indexes the neural network layer, with m representing the exponential operation, and $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$, with $N$ being the maximum number of neurons across all layers, $C_{s}$ representing a constant, and $K_{s}=\max\left\|\mathbf{W}^{(m)}\right\|_{2}$.
-This property shows that the output norm can grow exponentially with the number of layers $m$, controlled by the Lipschitz constant of the activation function and the bound on the weight matrices.
+This property shows that the output norm can grow exponentially with the number of layers $m$, controlled by the Lipschitz constant of the activation function and bounded by the max entry value of weight matrices.
 Theorem 1: Assume the weights in each layer are i.i.d.
 sub-Gaussian random variables with variance $\sigma_{w}^{2}$, and the biases are with variance $\sigma_{b}^{2}$.
-Furthermore, let the activation function $\phi$ be Lipschitz continuous with constant $L$.
-Then, the variance of the output of the final layer, represented by $V(f_{\theta}(x))$, can be bounded by:
-$\small\begin{split}V(f_{\theta}(x))\leq&\|x\|_{2}^{2}(L^{2}N)^{m}\prod_{m=1}^{M_{l}}(\sigma^{(m)}_{w})^{2}+L^{2}N\left(\sigma_{b}^{(M_{l})}\right)^{2}+\\ &L^{2}\sum_{m=1}^{M-1}\left\{N\left(\sigma_{b}^{(m)}\right)^{2}\prod_{l=m+1}^{M}\left[L^{2}N\left(\sigma^{(m)}_{w}\right)^{2}\right]\right\},\end{split}$ (16)
+Also, let the activation function $\phi$ be Lipschitz continuous with constant $L$.
+Then, the variance of the output of the final layer, represented by $\mathbb{V}(f_{\theta}(x))$, can be bounded by:
+$\small\begin{split}\mathbb{V}(f_{\theta}(x))\leq&\|x\|_{2}^{2}(L^{2}N)^{m}\prod_{m=1}^{M_{l}}(\sigma^{(m)}_{w})^{2}+L^{2}N\left(\sigma_{b}^{(M_{l})}\right)^{2}+\\ &L^{2}\sum_{m=1}^{M-1}\left\{N\left(\sigma_{b}^{(m)}\right)^{2}\prod_{l=m+1}^{M}\left[L^{2}N\left(\sigma^{(m)}_{w}\right)^{2}\right]\right\},\end{split}$ (16)
 where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant.
-This inequality shows that the output variance grows with the depth $M$ of the network and the variances of the weights and biases.
-Property 1 and Theorem 1 show that the increase of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances and vice-versa.
+This inequality shows that the output variance grows with the depth of the network, as well as impacted by the variances of the weights / bias.
+Property 1 and Theorem 1 show that the increment of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances.
 As shown in Fig.
@@
 Together, these factors can drive the weights to unstable values, reducing the model’s ability to generalize effectively.
-On the other hand, when $\mathbf{W}\to 0$, the output $\mathbf{y}$ will tend to a fixed point determined predominantly by the network’s bias terms, thus becoming largely independent of the input $\mathbf{x}$.
-Hence, if the magnitude of $\mathbf{W}$ is small, but not identically zero, then the range of outputs for $\mathbf{y}$ will be limited.
-This reduces the sensitivity of $\mathbf{y}$ to changes in $\mathbf{x}$, effectively shrinking the output space (similar to the L1 or L2 regularization).
+On the other hand, when $\mathbf{W}\to 0$, the output $\mathbf{y}$ will tend to a fixed point determined predominantly by the network’s bias terms, becoming largely independent of the input $\mathbf{x}$.
+Hence, if the magnitude of $\mathbf{W}$ is small, but not identically zero, then the range of outputs for $\mathbf{y}$ will be constrained.
+This reduces the sensitivity of $\mathbf{y}$ to the changes of $\mathbf{x}$, effectively shrinking the output space.
+In this case, model-merging acts similar to the L1 or L2 regularization.
 Such behavior can enhance the robustness of the model, as smaller weight magnitudes reduce the model’s response to minor perturbations in input $\mathbf{x}$.
-However, this comes with a trade-off: the expressive capacity of the model diminishes as well, potentially leading to poor performance on complex tasks.
+But this comes with a trade-off: the expressive capacity of models may diminish as well, potentially leading to compromised performance on complex tasks.
 {mdframed}
-[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec.
-6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging.
-This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness.
-In this case, model merging is similar to regularization.
-This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements.
-For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes may improve performance.
-However, for tasks requiring detailed pattern recognition, lower weight magnitudes may hinder the network’s ability to capture critical pattern distinctions in the data.
+[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analyses and experimental study in Sec.
+6.3, model-merging is more robust to the changes in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging.
+This is because, generally, the small magnitudes of weights enhances robustness by minimizing the network’s sensitivity towards input variations, but somehow reduces the expressiveness from ensembling all models.
+From this perspective, model-merging acts similarly as regularization.
+This balance between robustness and expressiveness is crucial and should be carefully considered based on the specific task requirements.
+For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes can improve performance.
 6 Experiments
-The experimental settings can be found in the supplementary materials.
+The experimental settings and details can be found in the supplementary materials.
 6.1 The Patterns Contained in Model Weights
@@
 As shown in Fig.
-1(a,b), the visualization of the linear classifier on Cifar100 clearly shows the patterns of the visual class learned, which are strongly correlated with the average of all images within each class.
-The template of each class acts as a merged collection of representative images of the respective class.
-Interestingly, if we merge the weights of the first 5 classes in Fig.
-1 with the last 5 classes, it gives us the results of Fig.
+1(a,b), the visualization of the linear classifier on Cifar100 explicitly shows the patterns per classifier learned, which are strongly correlated with the average of images per class.
+The patterns learned per class acts as a fused representative image.
+Interestingly, if we merge the weights of the first 5 classes with the last 5 classes in Fig.
+1, it gives us the results of Fig.
 2.
-As we analyzed before, the weight-averaged model-merging acts as the Mixup augmentation, which is a linear combination of different matrices, in the weight space.
+As mentioned, the weight-averaged model-merging acts as a Mixup operation, which is a linear combination of different matrices, in the weight space.
 For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle.
-We can also observe that after merging, each element within the combined template noticeably fades.
+We can also observe that, after merging, each element within the combined template noticeably fades.
 This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec.
@@
 Fig.
-6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes.
-For example, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visualization.
+6 visualizes the image average per class and its corresponding weight for Tiny ImageNet classes.
+For instance, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can see the shape of the cracker from its weight visualization.
 However, they are not as clear as for Cifar100.
-This may be due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear.
-This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100.
+This may be due to the relatively large dataset size, as the average of all images in a class itself is also vaguer.
+This compromised clarity can also be caused by the limited expressiveness of single linear classifiers and there are more complicated patterns present in Tiny ImageNet than Cifar100.
 Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [9] in Fig.
@@
 3.1.
-The visualization for every class in Cifar100 and Tiny ImageNet is put in the supplementary materials.
+The visualizations for all classes within Cifar100 and Tiny ImageNet are put in the supplementary materials.
 6.2 Weights vs. Features Averaging
@@
 As shown in Fig.
-3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst.
-We can also discover that greedy soups are generally better than uniform soups due to their mechanism to greedily select the best performing model to merge.
-It ensures that a good result can be reached by only incorporating the first model.
-The greedy ensemble on both logits and features also produces results that are comparable to greedy soups.
-In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensemble.
-Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups.
-This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs).
-CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for model merging.
-On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model merging.
-ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
+3 for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials).
+For abbreviations, “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively.
+Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures [9, 10, 24, 26, 5, 29].
+However, uniform soups perform the worst.
+We can also discover that greedy soups generally perform better than uniform soups due to their mechanism to greedily select the best performing model to merge.
+In fact, it guarantees that reasonably promising results can be achieved by only incorporating the first model (if the later-on merged model performance is not ideal).
+The greedy ensemble on both logits and features produces results that are comparable to greedy soups.
+In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensembling.
+Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models into.
+This happens may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs).
+CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for merging.
+On the contrary, ViTs have global attention mechanisms, where the templates represent attention and correlations between tokens, making them sensitive for small changes.
+ViT models also show similar deterioration on other data (as shown by the DeiT on TinyImageNet and it can be found in the supplementary materials).
 Figure 8: (a) For Sec.
@@
 Magnification Factor in terms of different model numbers on Cifar100 dataset.
-We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (as a function of the number of merged models and different model architectures, but excluding ViT models as their gaps are too large) in Fig.
-8.
-We can see that the more models in the merging process, the larger the gaps between logits ensemble and uniform soups.
+We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (excluding ViT models as the gaps are too large, so in different scales) in Fig.
+8(a).
+We can see that the more models in the merging process, the larger the gap between logits ensemble and uniform soups.
 The gaps are different according to different models, but generally, there is a downtrend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm).
-This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities.
-Batch normalization can break this downtrend.
-We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that greedy soups can alleviate this marginal effect.
+This may be caused by the performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities.
+But batch normalization can break this trend.
+We have also examined the gaps between logits ensemble and greedy soups in the supplementary materials, showing that greedy soups can alleviate this marginal effect.
 6.2.2 Observations over Different Datasets
@@
 ChestXRay 80.66 79.29 80.43 81.77 81.07 80.43 80.43
-For the dataset-wise examination, we perform model merging and ensemble on 7 prevalent datasets with different data sizes.
+For the dataset-wise examination, we perform model-merging and ensembling on 7 prevalent datasets with different data sizes (TinyImageNet and ChestXRay are relatively large, and both with more than 100,000 images in the datasets).
 From Tab.
-1, we can observe ensemble still generally performs better than merging.
-Different datasets have different performances and this is due to the characteristics of each dataset.
-Similarly to the model-wise comparison, with more models in the merging (2-7 models merging results in the supplementary materials), the better the performance for both merging and ensembling.
-The model ensembling is more stable than model merging in terms of the best and worst performance on different data.
-For example, on the PathMNIST [37, 38] dataset, uniform soup merging does not work really well compared to the relative performance on the Cifar10/Cifar100 or Tiny ImageNet datasets (that only achieves 40.85% accuracy, which is largely behind model ensemble methods).
-A similar situation happens on Celeba [15] dataset: the uniformly average model only achieves 53.19%.
-Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model.
-These results may be caused by the contradictions between the two models’ weights.
+1, we can observe that the ensembling still generally performs better than merging.
+The models on different datasets have different performances due to the dataset characteristics.
+Similarly to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling.
+The ensembling is more stable than merging in terms of the best and worst performance on different datasets.
+For example, on the PathMNIST [39, 40], uniform soup does not work very well (32.42% accuracy, which largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets.
+A similar situation shows on Celeba [17]: the uniformly average model only achieves 53.19%.
+Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering the first model.
+These results may be caused by the contradictions between model weights.
 6.2.3 Merging Models trained over different Tasks
 Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods.
-Model 1 is the VGG19 model trained on the Cifar100 dataset, while Model 2 denotes the VGG19 model trained on each different dataset with different tasks.
-Feature ensemble is better than logits ensemble.
-Datasets Size Model1 Model2 Uniform Soups Ens Lgt Ens Features
-Cifar10 60,000 70.37 5.84 49.63 69.93 69.93
-PathMNIST 107,180 70.37 1.29 24.25 69.88 70.00
-DermaMNIST 10,015 70.37 5.77 60.37 69.06 69.36
-Celeba 202,599 70.37 1.03 49.33 69.48 69.61
-TinyImageNet 100,000 70.37 18.16 61.02 58.40 58.52
-ChestXRay 112,120 70.37 15.18 61.38 68.18 68.33
-We also explore model merging between different well-trained models on different tasks, as shown in Tab.
+Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
+E.g.
+Model 2 trained on Cifar10 receives 5.84% accuracy on Cifar100 dataset.
+Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
+Cifar10 60,000 5.84 49.63 69.93 69.93
+PathMNIST 107,180 1.29 24.25 69.88 70.00
+DermaMNIST 10,015 5.77 60.37 69.06 69.36
+Celeba 202,599 1.03 49.33 69.48 69.61
+TinyImageNet 100,000 18.16 61.02 58.40 58.52
+ChestXRay 112,120 15.18 61.38 68.18 68.33
+We also explore model-merging between different well-trained models on different tasks, as shown in Tab.
 2.
-Generally, merging within Tiny ImageNet (where Model 2 was trained on) has the best performance of 61.02% accuracy, which is better than the ensemble on logits (58.40% accuracy).
-This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images.
-Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results do not perform better than the model ensemble, as it is too small.
-On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement.
-This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data was learned.
-However, the performance on Celeba is not good.
-This may be caused by the Celeba dataset only containing human faces, which can be seen as a fine-grained classification task.
-Due to the uniformity of the data format within Celeba, it contradicts the diverse and rich general computer vision features learned by the model weights on Cifar100.
+Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
+Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy).
+It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features.
+Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling.
+Maybe because it is too small.
+In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along.
+It illustrates that the dataset size also matters.
+The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).
 6.3 Model Predictions and Weight Magnitudes
 […]
-Specifically, we visualize the merging of 10 models (merging of 2 to 7 models can be found in the supplementary materials) for model-merging and ensemble towards different magnitude factors, as shown in Fig.
+Moreover, we visualize 10 models (merging of 2 to 7 models can be found in the supplementary materials) merging and ensembling towards different magnitude factors, as shown in Fig.
 8(b).
-In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with weight magnification.
-From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features.
-But surprisingly, feature ensembles are more robust than logits ensembles across different configurations.
-Feature ensemble models can even outperform model merging when weight magnification is not large (e.g., ×$\times 50$).
-This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances.
-The “Ens Features” model uses the merge of the last fully connected (FC) layers from all merged models for classification.
+In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with different magnification.
+From the trends in the figures, we can observe the merged models are more robust than ensembled ones (with the factor grows, the line denoting ensembling goes down quickly), including ensemble on logits and features.
+But interestingly, feature ensembling is more robust than logits ensembling across different configurations.
+Feature ensembling can even outperform model-merging when weight magnification is not large (e.g., ×$\times 50$).
+This shows the feature ensembling can have some resistance towards the change of weight magnitudes.
+The “Ens Features” performs ensemble on multiple model features, while merging all last fully connected (FC) layers to get rid of the influence of FC layer.
 In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model.
 We can observe the “Ens Features” and “Ens Features*” models perform similarly.
-We further visualize the gaps between weight-averaged model-merging and ensemble in Fig.
+It shows the low sensitivity of the merging of last layers.
+We further visualize the gaps between model-merging and ensembling in Fig.
 8(c).
-When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ for Cifar100).
-The same experiments to visualize the merging of 2 to 10 models for Cifar10 with different magnitude factors, as well as the gaps between weight-averaged model-merging and ensemble, are shown in the supplementary materials.
-Also, a similar phenomenon is shown on Cifar10, but with different peaks of gaps w.r.t.
-the magnitudes.
+When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ is the peak for Cifar100).
+The experiments of merging/ensembling 2 to 10 models with different magnitude factors, as well as on different datasets and the gaps between merging and ensembling, are shown in the supplementary materials.
+A similar phenomenon is shown on different datasets, but with different peaks.
 Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab.
 3.
-From the table, we can observe that model merging can still mitigate the weight magnifying problem to a certain extent.
-It can also outperform logit ensembles, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g.
-4.81% to 6.86% accuracy for 10 models merging).
-This may be caused by the complexity of the Tiny ImageNet task and as we discussed in the Sec.
-5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization).
-It will mitigate the weight magnitude changing issues, but with the cost of losing model expressiveness.
+From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent.
+It can also outperform logit ensembling, and cannot outperform feature ensembling.
+But generally it cannot work very well as on Cifar100.
+It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec.
+5.3, the model-merging can work to some extent as it can act similarly as regularization.
+It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
 Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$.
-This approach works for smaller data (Cifar10/Cifar100) as it plays a regularization role.
+This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect.
 However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
@@
 In this paper, we thoroughly examined the weight-averaged model-merging approach through the lens of template matching by visualizing weight patterns across various datasets.
-Our comparison of weight and feature averaging within ensemble and merging methods revealed notable differences in their effects on different model architectures and datasets, offering insights into when and why each approach is beneficial.
-Furthermore, we demonstrate that model merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization.
+Our comparison of weight and feature averaging within ensembling and merging methods revealed notable differences in their effects on different architectures and datasets, offering insights into when and why each approach is beneficial.
+Furthermore, we demonstrate that model-merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization.
 These findings clarify the interpretability and key aspects of model-merging, providing a deeper understanding of its application across diverse contexts.
-We anticipate that these insights will stimulate further research into model merging strategies and support the development of more interpretable and robust model-averaging methodologies.
+We anticipate that these insights will stimulate further research into model-merging strategies and support the development of more interpretable and robust model-averaging methodologies.
 Supplementary Material
@@
 The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1.
-While visualization, the weights are wrapped back to the image size.
-In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging.
+While visualization, the weights are wrapped back to the image size as described in the main paper (i.e.
+100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
+In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [35] for weight-averaged model-merging, which is the focus of our paper.
 We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances.
@@
 4.
-Specially, ChestXRay dataset has 14 labels and each of them is a binary classification task.
+Specially, ChestXRay-14 is a chest X-ray imaging dataset designed for multi-label classification.
+It consists of 112,120 frontal chest X-ray images from 30,805 unique patients.
+The dataset includes annotations for 14 different lung disease categories.
+We use the official dataset split for training and evaluation.
 We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay.
-For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer).
-Besides these two model ensemble techniques, whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts).
+For model ensembling, we consider to ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer).
+Besides these two model ensembling techniques, whenever a model is considered to be merged into a greedy soup model, we also ensemble it with either we called the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts).
 In order to exclude the impact of adopting the last layer from any single model, the “Ens Features” model uses the merging of the last fully connected (FC) layers from all merged models for classification.
 For the experiments of model predictions and weight magnitudes, we conduct experiments on Cifar10, Cifar100 and Tiny ImageNet datasets.
-To increase the magnitude and the variance of model weights, we magnify the model weights by a constant factor.
-Formally, when each element in a matrix $A$ is scaled by a constant factor $c$, the variance of the resulting matrix elements increases proportionally to the square of this factor.
+To increase the magnitude and the variance of model weights, we magnify the model weights by a magnitude factor.
+Formally, when each element in a matrix $A$ is scaled by a factor $c$, the variance of the resulting matrix elements increases proportionally to the square of this factor.
 For example, in the case where $c=100$, the variance of the elements of matrix $A$ becomes $100^{2}\times\text{Var}(A)=10,000\times\text{Var}(A)$.
 For all experiments in the paper, to keep fair comparisons, we adopt the official splits for each dataset.
-Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions (learning rates, step strategies and optimizers, etc).
+If they provide publicly available dataset with Python interface, we directly call the interface to import the data.
+Also, we uniformly train 10 models and perform the model-merging and ensembling under exactly same conditions (learning rates, step strategies and optimizers, model numbers, etc).
+9 Theoretical Proofs
+For Eq. 14 in the paper, we aim to prove that:
+$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\max(\sigma_{1}^{2},\sigma_{2}^{2})$
+Assume without loss of generality that $\sigma_{1}^{2}\leq\sigma_{2}^{2}$ (the argument is symmetric if $\sigma_{2}^{2}\leq\sigma_{1}^{2}$).
+Given that $\max(\sigma_{1}^{2},\sigma_{2}^{2})=\sigma_{2}^{2}$, the inequality becomes:
+$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\sigma_{2}^{2}$
+Now, observe the following:
+$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\frac{1}{4}(2\sigma_{2}^{2})=\frac{\sigma_{2}^{2}}{2}$
+Since $\frac{\sigma_{2}^{2}}{2}\leq\sigma_{2}^{2}$, we conclude that:
+$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\sigma_{2}^{2}=\max(\sigma_{1}^{2},\sigma_{2}^{2})$
+Thus, the inequality holds true.
+For the Property 1 and Theorem 1, according to [34], we have:
+Lemma 1: Given a fully-connected network, its weight matrix $\mathbf{W}^{(m)}$ satisfies, for any $\tau>0$,
+$\small\norm{\mathbf{W}^{(m)}}_{2}\leq\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau),$
+with a probability of at least $1-2e^{-\tau^{2}}$, where $C_{s}$ is a universal constant and $K_{s}=\max_{i}\|\mathbf{W}_{i}^{(m)}\|_{2}$.
+For Property 1:
+Proof.
+We denote $s_{1}(A)$ as the maximum singular value of $A$ and set $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$ as the upper bound of the maximum singular value of all weight matrices by Lemma 1.
+By Lipschiz property of the activation function $\phi$ and $\norm{A\mathbf{x}}=s_{1}(A)\norm{\mathbf{x}}$,
+$\norm{\mathbf{y}^{(m)}}_{2}=$ $\norm{\phi(\mathbf{W}^{(m)}\mathbf{y}_{k}^{(m-1)})}_{2}$
+$=$ $\norm{\phi(\mathbf{W}^{(m)}\mathbf{y}_{k}^{(m-1)})-\phi(0)}_{2}$
+$\leq$ $L\norm{\mathbf{W}^{(m)}\mathbf{y}^{(m-1)}-0}_{2}$
+$=$ $L\norm{\mathbf{W}^{(m)}\mathbf{y}^{(m-1)}}_{2}$
+$=$ $Ls_{1}(\mathbf{W}^{(m)})\norm{\mathbf{y}^{(m-1)}}_{2}$
+$\leq$ $L\lambda\norm{\mathbf{y}^{(m-1)}}_{2}$
+$\leq$ $(L\lambda)^{2}\norm{\mathbf{y}^{(m-2)}}_{2}$
+$\cdots$
+$\leq$ $(L\lambda)^{m}\norm{\mathbf{x}}_{2}.$
+Then we calculate the probability that this inequality holds by Lemma 1,
+$\mathbb{P}\left(\norm{\mathbf{y}^{(m)}}_{2}\leq(L\lambda)^{m}\norm{\mathbf{x}}_{2}\right)$
+$=$ $\prod_{k=1}^{m}\mathbb{P}\left(\norm{\mathbf{W}^{(k)}\mathbf{y}^{(k-1)}}_{2}\leq s_{1}(\mathbf{W}^{(k)})\norm{\mathbf{y}^{(k-1)}}_{2}\right)$
+$\geq$ $(1-2e^{-\tau^{2}})^{m}.$
+For Theorem 1:
+Proof.
+For the first hidden layer, we have a linear transformation for one neuron:
+$\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i}.$
+Given that all weights and biases are i.i.d, we get:
+$\mathbb{V}(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})$ $=\mathbb{V}\left(\sum^{N^{(1)}}_{j=1}\mathbf{W}^{(1)}_{i,j}x_{j}+b^{(1)}_{i}\right)$
+$=\mathbb{V}\left(\sum^{N^{(1)}}_{j=1}\mathbf{W}^{(1)}_{i,j}x_{j}\right)+\mathbb{V}(b^{(1)}_{i})$
+$=\sum^{N^{(1)}}_{j=1}x_{j}^{2}\mathbb{V}(\mathbf{W}^{(1)}_{i,j})+\mathbb{V}(b^{(1)}_{i})$
+$=(\sigma^{(1)}_{w})^{2}\sum^{N^{(1)}}_{j=1}x_{j}^{2}+(\sigma^{(1)}_{b})^{2}$
+$=(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}.$
+By the Lipschitz property of the activation function, we have:
+$f^{(1)}_{i}=\phi(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})\leq L(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i}).$
+Then,
+$\mathbb{V}(f^{(1)}_{i})$ $\leq\mathbb{V}[L(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})]$
+$\leq L^{2}\mathbb{V}(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})$
+$\leq L^{2}\left[(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}\right].$
+Hence,
+$\mathbb{V}(f^{(1)})$ $=\sum^{N}_{j=1}\mathbb{V}(f^{(1)}_{i})$
+$\leq L^{2}N\left[(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}\right].$
+Similarly, for other layers:
+$\mathbb{V}(f^{(m)})\leq L^{2}N^{(m-1)}\left[(\sigma^{(m)}_{w})^{2}\|f^{\prime(m-1)}\|_{2}^{2}+(\sigma^{(m)}_{b})^{2}\right].$
+This result follows by iteratively applying the inequalities, concluding the proof. ∎
 10 More Visualization of Model Weight Patterns
@@
 10 within each class (all 100 classes) of Cifar100 dataset.
-We can easily and clearly observe the strong correlation between the average of images in a class and the learned weights.
+We can clearly observe the strong correlation between the average of images per class and its learned weights.
 Figure 11: Merge the linear classifiers for the first 50 classes with the last 50 classes in Cifar100 dataset.
@@
 For ResNet50 kernel visualization, we can observe differences or similarities between each learned patterns.
-For instance, Kernel 5 and kernel 9 are similar, which are responsible for detecting similar features and thus yield a higher measurement (e.g.
-inner product) between the two kernels; while Kernel 9 is very different with Kernel 45.The patterns inside ResNet50 is easier to be observed than VGG19, as ResNet50 has larger kernel size for the 1st layer.
+For instance, Kernel 5 and kernel 9 are similar in terms of textures and grayscale colors, which are responsible for detecting similar features and thus yield a similar activations (e.g.
+inner product) when interact with inputs; while Kernel 9 is very different with Kernel 45.
+The patterns inside ResNet50 is more human-friendly to be observed than VGG19, as ResNet50 models have larger kernel sizes.
 Figure 14: Average of all images within each class of TinyImageNet dataset (the first 100 classes).
@@
 Figure 17: Linear classifiers visualization for each class of TinyImageNet dataset (the last 100 classes).
-For Tiny ImageNet, we have Fig.
+In terms of Tiny ImageNet, we have Fig.
 14 and Fig.
-16 for the average of all images within 200 classes in the dataset; also, Fig.
+16 for the average of all images within 200 classes; also, Fig.
 15 and Fig.
 17 for the linear classifier visualization of all 200 classes in the dataset.
-We can observe the correlations between averaged images and the learned patterns.
+We can observe the correlations between averaged images and the learned patterns show a similar phenomenon.
 11 More Experimental Results of Averaging on Weights V.S. Averaging on Features
@@
 22 shows the performance gaps between ensembling on logits with greedy soups.
-From it, we can observe similar phenomenon with the gaps between ensembling on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be.
-But differently, the greedy soups break the going down trend with uniform soups as described in the main paper.
+From it, we can observe the similar phenomenon with the gaps of uniform soups, where the gaps grow when more models merged.
+But differently, the greedy soups break the going down trend as it can guarantee to maintain a relatively high accuracy level by only include the first model.
 Table 5: Performance Comparison for VGG19 with 2 Models Across Different Datasets.
@@
 10.
-We can also see the more models to be merged, the poorer results from uniform soups can get.
-Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
+We can also see the more models are merged, the worse results from uniform soups can get.
+Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed.
+It demonstrates the model-merging between 2 or more ViT models will not contribute to the performance.
+Thus, it shows the ViT models are not suitable for model-merging in the tested scenarios.
 12 More Experimental Results of Model Magnitude Changes
@@
 Figure 26: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
-The 2 models, 3 models, 5 models and 7 models accuracy vs. magnification factor on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
+The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
 Figure 27: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset.
@@
 Moreover, we also visualize the uniform soup and logits ensemble accuracy vs.
-magnification factor across different model numbers on the Cifar10 dataset in Fig.
+different magnification factors across different model numbers on the Cifar10 dataset in Fig.
 32.
-We can observe similar phenomenons.
-However, on the Cifar10, the gaps have different a peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).
+We can observe similar phenomenons as discussed in the apepr.
+However, on the Cifar10, the gaps have different peak points than Cifar100, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v4

@@
 However, the underlying mechanisms that explain its effectiveness remain largely unexplored.
-In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging works: (1) we examine the intrinsic patterns captured by the learning of the model weights, through the visualizations of their patterns on several datasets, showing that these weights often encode structured and interpretable patterns and that is the essential why model-merging can work; (2) we mathematically and empirically investigate model ensemble merging strategies based on averaging on weights versus averaging on features, providing detailed analyses across diverse architectures and datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales.
-Our findings shed light on the “black box” of weight-averaged model-merging, offering valuable insights and practical recommendations that advance the model-merging process.
+In this paper, we investigate this technique from three novel perspectives to empirically provide deeper insights into why and how weight-averaged model-merging Wortsman et al.
+(2022) works: (1) we examine the intrinsic patterns captured by the learning of the model weights, and we are the first to connect that these weights encode structured with why weight-averaged model merging can work; (2) we investigate averaging on weights versus averaging on features, providing analyses from the view of diverse architecture comparisons on multiple datasets; and (3) we explore the impact on model-merging prediction stability in terms of changing the parameter magnitude, revealing insights into the way of weight averaging works as regularization by showing the robustness across different parameter scales.
 The code is available at https://github.com/billhhh/Rethink-Merge.
@@
 Model-merging can combine multiple independently trained models into a single, effective model without introducing extra inference computations.
-This family of approaches has been applied in contexts such as natural language processing [38], computer vision [16], and their sub-areas (e.g.
-federated learning [32], distillation [13], adversarial learning [44, 2] and large language model [22, 45]), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment.
+This family of approaches has been applied in contexts such as natural language processing Yang et al.
+(2024), computer vision Li et al.
+(2023), and their sub-areas (e.g.
+federated learning Wang et al.
+(2020), distillation Khanuja et al.
+(2021), adversarial learning Zhang et al.
+(2024); Croce et al.
+(2023) and in large language models Rame et al.
+(2024); Zhou et al.
+(2024)), where the ability to merge models effectively can lead to largely improved models’ performance, efficiency, and flexibility in deployment.
 Despite the practical importance of model-merging, there has been limited research on the investigation of the underlying factors that make it effective.
-This research gap motivates the need to innovatively explore the underlying principles of why and how the family of weight-averaged model-merging methods can work.
+This research gap motivates the need to explore the underlying principles of why and how the family of weight-averaged model-merging methods can work.
 In this paper, we bridge this gap by providing a comprehensive analysis of weight-averaged model-merging from the following three key perspectives:
-  • An examination of the patterns contained in the model weights through the visualization of the patterns present in both linear layers as classifiers or inside deep learning models used on multiple datasets.
-Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be seen as a linear vector combination, which contributes to the success of merging models via weight operations;
-  • An investigation of averaging on weights versus averaging on features strategies allows us to evaluate the relative effectiveness and limitations of model-merging and traditional model ensembles across multiple architectures and datasets, uncovering distinct benefits and behaviors;
-  • An exploration of the robustness of model-merging predictions with respect to different weight magnitudes/variances, which provides insights into the model-merging works as a type of model regularization.
-Based on our viewpoints and findings, we conclude with useful proposals to provide a good reference for the community.
+  • We are the first to connect the patterns contained in the model weights to model-merging.
+It is conducted through the visualization of the patterns present in both linear classifiers or inside deep models on multiple datasets.
+Results reveal that clear and structured patterns emerge within the weights, suggesting that weight averaging can be viewed as a linear combination of two weight vectors;
+  • We present a structured comparison between weight averaging and feature averaging strategies, focusing on their implications for model merging and ensembling.
+Our analysis across diverse models and datasets reveals consistent behavioral differences, offering a more nuanced understanding of model merging and ensembling behavior with different architectures;
+  • We systematically analyze how the robustness of model-merging predictions is affected by differences in weight magnitudes and variances.
+Our findings reveal that model merging implicitly acts as a form of regularization, favoring solutions with smoother or more stable parameter configurations.
+This connection provides a theoretical perspective on when and why model merging improves generalization.
+The intention of the paper: Instead of providing a thorough evaluation of existing works, e.g.
+Ainsworth et al.
+(2022); Yadav et al.
+(2024); Jang et al.
+(2024); Singh and Jaggi (2020), as a survey paper, the main focus of this paper is to offer a systematic analysis of why and how weight-averaged model-merging  Wortsman et al.
+(2022) can work from the abovementioned three interesting viewpoints.
+Based on our viewpoints and findings, we also conclude with useful proposals to provide a good reference for the community.
 2 Related Work
@@
 (b) Linear classifiers visualization.
-Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 [14] dataset (b).
+Figure 1: Average of all images for each class (a) and the linear classifier visualization for each class on Cifar100 Krizhevsky (2009) dataset (b).
 We train a single linear classifier for all classes then wrap it back to the size of input images.
@@
 It interprets the reason why the model-merging techniques can work as they essentially operate the templates contained in model weights.
-In recent years, model-merging [28, 23, 3, 21, 25, 37] has attracted growing interest.
+In recent years, model-merging Sung et al.
+(2023); Sanjeev et al.
+(2024); Daheim et al.
+(2023); Navon et al.
+(2023); Singh and Jaggi (2020); Yang et al.
+(2023a); Jang et al.
+(2024); Almakky et al.
+(2024); Singh et al.
+(2024) has attracted growing interest.
 This approach has led to a range of methods focused on merging weights from independently trained models to enhance model performance or generalization.
 For example, Wortsman et al.
-[35] proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference.
+Wortsman et al.
+(2022) proposed an approach of averaging, or greedily averaging, the parameters from various models to yield “Model Soups” that show more accurate performance than the original models used to produce the “soup” but without introducing more computations into inference.
 Further, Yadav et al.
-proposed Ties-Merging [36] based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models.
-Some other approaches [1, 6, 12] address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins.
-Zipit [27] takes one step ahead from Git Re-Basin [1] to align weights from inter-model and intra-model based on the alignment of features between merging models.
+proposed Ties-Merging Yadav et al.
+(2024) based on parameter averaging by considering conflicts between the model weights, allowing for smooth parameter merging across models.
+Some other approaches Ainsworth et al.
+(2022); Entezari et al.
+(2021); Jordan et al.
+(2022) address the problem from a different perspective, focusing on taking weight merging operations via finding a good way to interpolate on the loss basins.
+Zipit Stoica et al.
+(2023) takes one step ahead from Git Re-Basin Ainsworth et al.
+(2022) to align weights from inter-model and intra-model based on the alignment of features between merging models.
 Yu et al.
-[41] proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
+Yu et al.
+(2024) proposed a DARE model to combine models trained on domain-specific data to create a domain-agnostic model-merging that generalizes well across diverse distributions.
 This approach ensures that domain knowledge is preserved while enhancing the model’s ability to generalize across different tasks.
-Beyond model-merging, other techniques, such as mode connectivity path ensembles [8], analyze pathways in the parameter space between models.
+Beyond model-merging, other techniques, such as mode connectivity path ensembles Garipov et al.
+(2018), analyze pathways in the parameter space between models.
 By identifying optimal modes along these paths, these models help improve the performance after merging.
-Fisher Merging [19] selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
-Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited.
-This gap in understanding serves as the motivation for our paper.
-To focus on a specific scope, we aim to offer a thorough analysis from mathematical and empirical perspectives of weight-averaged model-merging different from the study in [35], which serves as the foundation work for the area.
+Fisher Merging Matena and Raffel (2022) selects parameters that approximately maximize the joint likelihood of the models’ parameter posteriors.
+Despite these advances, research into the underlying mechanisms that make model-merging effective remains limited. This gap in understanding serves as the motivation for our paper.
 3 The Patterns Contained in Model Weights
-In this section, we will explain the reason why does the weight-averaged model-merging work via the analysis of their weights from the perspective of template matching (an example is shown Fig.
-1 and Fig.
-2 for a linear model; for nonlinear models, visualizations can be found in the supplementary materials).
-Specifically, we explain the reason that the weights can be viewed as templates learned from the data and build connections with an inner product which is a widely explored operation in neural networks.
-By treating the weights as templates, it is understandable that operations in weight space as efforts to manipulate feature representations.
 3.1 Connections Between Inner Product and Template Matching
 The weights of the model can be interpreted as templates or prototypes for the respective class activations and it has a deep connection with machine learning.
-A couple of decades ago, Adaboost [7] contained heuristically designed patterns such as HAAR features [31], which are simple hand-crafted filters that capture edges / lines / other low-level image features through weighted sums and differences in pixel intensities.
+A couple of decades ago, Adaboost Freund et al.
+(1999) contained heuristically designed patterns such as HAAR features Viola and Jones (2001), which are simple hand-crafted filters that capture edges/lines/other low-level image features through weighted sums and differences in pixel intensities.
 During training, Adaboost iteratively adjusts the weights of weak classifiers to improve their performance on misclassified examples.
-In a linear classification, once the weights are learned, we wrap it back to have the same size with the input images.
+In a linear classification, once the weight matrix is learned, we wrap it back to have the same size as the input images.
 Take Cifar100 for instance, each image within the dataset is 32$\times$32$\times$3, if we flatten it and it will become a 3072$\times$1 vector; the weight matrix size is 100$\times$3072 without considering bias; in this case, each row of the weight matrix corresponds to each class to be classified and thus each row vector (1$\times$3072) can be wrapped back to 32$\times$32$\times$3 with the same size of the original images.
@@
 For each class, which weight vector yields the highest activation is selected as the predicted class of that vector, indicating that the “template” aligns (or similar to) the input image most closely.
-This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g.
-in the lower layer, low level features, e.g.
-edge / line features, are tend to be learned).
-The averaging operation in the weight space represents as Fig.
-2, which can be seen as the Mixup [42] operation in the weight space.
-Therefore, the linear combinations in the weight space has a similar effect as in the feature space.
-This also explains why recent research [4, 43, 11, 46] applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec [20].
+This phenomenon also holds for deeper networks, but differently from the linear classifier, weight templates in each layer are learned to detect different features that contribute the most towards the objectives (e.g., in the lower layer, low-level features, e.g., edge/line features, are tend to be learned).
+The averaging operation in the weight space is represented as Fig.
+2, which can be seen as the Mixup Zhang (2017) operation in the weight space.
+Therefore, the linear combinations in the weight space have a similar effect as in the feature space.
+This also explains why recent research Dekoninck et al.
+(2023); Zhang et al.
+(2023); Ilharco et al.
+(2022); Zhou et al.
+; Ortiz-Jimenez et al.
+(2023) applies “model arithmetic” on multiple models in the weight space can achieve the effect of feature arithmetic as Word2vec Mikolov (2013).
 Extending to model-merging, the combined model retains the diverse patterns learned from each ingredient model.
@@
 On the other hand, when $\theta\neq 0$, then $\cos\theta<1$, which means that the template and image features become misaligned.
-This mechanism also applies for the layers within deep learning models.
+This mechanism also applies to the layers within deep learning models.
 The visualization of deep models can be found in the supplementary materials.
@@
 4 Averaging on Weights vs. Averaging on Features
-Figure 3: The 10 model-merging/ensemble accuracy comparison between different architectures on Cifar100.
+Figure 3: (a) The 10 model-merging/ensemble accuracy comparison between different architectures on Cifar100.
 Comparison between averaging on Weights (Uniform Soups, Greedy Soups), averaging on features (Ens Logits, Ens Features, Greedy Ens Logits, Greedy Ens Features) and original model performance (Performance Average) on Cifar100 using many different deep learning architectures, including CNNs and ViTs.
@@
 6.2.
-Averaging on weights (model-merging) and Averaging on features (model ensembling) are two different ways to aggregate knowledge from multiple models without retraining.
-These methods have distinct properties and lead to different testing behaviors.
-In this section, we explore the differences between these two approaches and analyze their effects on the model effectiveness and efficiency, one example experiment is shown in Fig.
-3.
+(b) Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100.
+The values on the X and Y axis represent the accuracies.
+The experimental analysis is in Sec.
+6.3.
+In this section, we explore the differences between averaging on weights and features by analyzing their effects on the model effectiveness and efficiency, one example experiment is shown in Fig.
+3(a).
 4.1 Linear Model Scenarios
@@
 In such a case, the equivalency between $\overline{\mathbf{h}}$ and $\overline{\mathbf{h}}^{\prime}$ seen in the linear models may not hold for non-linear models.
-Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to large loss of information after the non-linear activation.
+Take ReLU for instance, if elements of $\mathbf{W}_{1}\mathbf{x}$ and $\mathbf{W}_{2}\mathbf{x}$ exhibit totally opposite activations (same absolute values but with different signs), where their average will tend towards zero, leading to a large loss of information after the non-linear activation.
 Therefore, averaging on weights tends to smooth out the contributions from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, which can be beneficial in certain scenarios by reducing variance or providing a regularizing effect.
@@
 5 The Model Predictions and Weight Magnitudes
-Figure 4: Scatter plot comparing the classification accuracies of model merge (uniform soup) and model ensemble (logits ensemble) over different weights magnitudes with VGG19 model on Cifar100.
-The values on the X and Y axis represent the accuracies.
-The experimental analysis is in Sec.
-6.3.
-In this section, we will examine how the weight-merging behaves against magnifying the candidate model weights by different magnitude factors. This helps us understand how robust model-merging works.
+In this section, we will examine how the weight-averged merging behaves when the candidate model weights are multiplied by different multiples (e.g., 90×, 100×, etc.), defined as magnitude factors.
+This helps us understand how robust model-merging works.
 5.1 Analysis of Weight Magnitude After Averaging
-We start from analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes.
+We start by analyzing the effect of averaging the weights of two independent matrices, denoted as $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, on their respective magnitudes.
 We define the maximum entry of both matrices as $M=\max(\mathbf{W}_{1},\mathbf{W}_{2})=\max_{i,j}(|(\mathbf{W}_{1})_{i,j}|,|(\mathbf{W}_{2})_{i,j}|)$.
@@
 We seek to understand the relationship between the max entry of the averaged weights $\overline{\mathbf{W}}$ and $M$.
-Mathematical Justification. Using the triangle inequality [30], we establish:
+Mathematical Justification. Using the triangle inequality Tversky and Gati (1982), we establish:
 $\small\|\overline{\mathbf{W}}\|=\left\|\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right\|\leq\frac{\|\mathbf{W}_{1}\|+\|\mathbf{W}_{2}\|}{2},$ (10)
@@
 10.
-It confirms that averaging the weights results in a new weight whose entry’s upper bound is equal to the maximum entry of the original weights.
+It confirms that averaging the weights results in a new weight whose entries’ upper bound is equal to the maximum entry of the original weights.
 Thus, the operation of averaging the weights $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ typically results in a reduction of the maximum weight magnitude.
@@
 5.3 How Weight Magnitudes and Variance Effects Model Outputs
-Figure 5: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of the decision boundary.
-We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  [34]:
+We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  Wang et al. (2024):
 Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ is the parameters.
@@
 where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant.
-This inequality shows that the output variance grows with the depth of the network, as well as impacted by the variances of the weights / bias.
+This inequality shows that the output variance grows with the depth of the network, and is impacted by the variances of the weights/bias.
 Property 1 and Theorem 1 show that the increment of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances.
 As shown in fig.
-5, consider a neural network with input $\mathbf{x}$ and output $\mathbf{y}$, where weights are denoted as $\mathbf{W}$.
+6, consider a neural network with input $\mathbf{x}$ and output $\mathbf{y}$, where weights are denoted as $\mathbf{W}$.
 When model weights have large magnitudes and high variance, predictions become unreliable.
@@
 6.3, model-merging is more robust to the changes in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging.
-This is because, generally, the small magnitudes of weights enhances robustness by minimizing the network’s sensitivity towards input variations, but somehow reduces the expressiveness from ensembling all models.
-From this perspective, model-merging acts similarly as regularization.
+This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but somehow reduce the expressiveness from ensembling all models.
+From this perspective, model-merging acts similarly to regularization.
 This balance between robustness and expressiveness is crucial and should be carefully considered based on the specific task requirements.
@@
 (b) Linear classifiers visualization.
-Figure 6: Average of all images for each class and linear classifiers visualization for each class in Tiny ImageNet [15] dataset.
-Figure 7: The examples of the first layer weights visualization of Resnet50 with nonlinearity inside the model.
+Figure 4: Average of all images for each class and linear classifiers visualization for each class in Tiny ImageNet Le and Yang (2015) dataset.
 As shown in Fig.
 1(a,b), the visualization of the linear classifier on Cifar100 explicitly shows the patterns per classifier learned, which are strongly correlated with the average of images per class.
-The patterns learned per class acts as a fused representative image.
+The patterns learned per class act as a fused representative image.
 Interestingly, if we merge the weights of the first 5 classes with the last 5 classes in fig.
@@
 We can also observe that, after merging, each element within the combined template noticeably fades.
-This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec.
-4.
+This phenomenon can also explain the limitation of expressiveness of the merged model as described in section 4.
 fig.
@@
 This compromised clarity can also be caused by the limited expressiveness of single linear classifiers and there are more complicated patterns present in Tiny ImageNet than Cifar100.
-Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [9] in Fig.
+Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model He et al.
+(2016) in fig.
 7 (VGG model weight visualization can be found in the supplementary materials).
@@
 As shown in Fig.
-3 for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials).
+3(a) for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials).
 For abbreviations, “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively.
-Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures [9, 10, 24, 26, 5, 29].
+Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures He et al.
+(2016); Huang et al.
+(2017); Simonyan (2014); Steiner et al.
+(2021); Dosovitskiy (2020); Touvron et al.
+(2021).
 However, uniform soups perform the worst.
@@
 In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensembling.
-Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models into.
-This happens may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs).
-CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for merging.
+Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models.
+It may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs).
+CNNs are more adapted to the learning of local patterns within different scales of images, making CNN architectures ideal for merging.
 On the contrary, ViTs have global attention mechanisms, where the templates represent attention and correlations between tokens, making them sensitive for small changes.
@@
 We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (excluding ViT models as the gaps are too large, so in different scales) in fig.
-8(a).
+5(a).
 We can see that the more models in the merging process, the larger the gap between logits ensemble and uniform soups.
@@
 Table 1: Performance Comparison for VGG19 with 10 Models Across Different Datasets.
-We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay [33, 18] dataset; while accuracy (0% to 100%) for the rest of datasets.
+We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay Wang et al.
+(2017); Ma et al.
+(2019) dataset; while accuracy (0% to 100%) for the rest of datasets.
 Datasets Perf Ave Uni Soups Grd Soups Ens Lgt Ens Fts Grd Ens Lgt Grd Ens Fts
@@
 The models on different datasets have different performances due to the dataset characteristics.
-Similarly to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling.
+Similar to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling.
 The ensembling is more stable than merging in terms of the best and worst performance on different datasets.
-For example, on the PathMNIST [39, 40], uniform soup does not work very well (32.42% accuracy, which largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets.
-A similar situation shows on Celeba [17]: the uniformly average model only achieves 53.19%.
+For example, on the PathMNIST Yang et al.
+(2021, 2023b), uniform soup does not work very well (32.42% accuracy, which is largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets.
+A similar situation shows on Celeba Liu et al.
+(2015): the uniformly average model only achieves 53.19%.
 Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering the first model.
 These results may be caused by the contradictions between model weights.
-6.2.3 Merging Models trained over different Tasks
-Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods.
-Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
-E.g.
-Model 2 trained on Cifar10 receives 5.84% accuracy on Cifar100 dataset.
-Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
-Cifar10 60,000 5.84 49.63 69.93 69.93
-PathMNIST 107,180 1.29 24.25 69.88 70.00
-DermaMNIST 10,015 5.77 60.37 69.06 69.36
-Celeba 202,599 1.03 49.33 69.48 69.61
-TinyImageNet 100,000 18.16 61.02 58.40 58.52
-ChestXRay 112,120 15.18 61.38 68.18 68.33
-We also explore model-merging between different well-trained models on different tasks, as shown in Tab.
-2.
-Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
-Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy).
-It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features.
-Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling.
-Maybe because it is too small.
-In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along.
-It illustrates that the dataset size also matters.
-The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).
 6.3 Model Predictions and Weight Magnitudes
@@
 Moreover, we visualize 10 models (merging of 2 to 7 models can be found in the supplementary materials) merging and ensembling towards different magnitude factors, as shown in fig.
-8(b).
+5(b).
 In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with different magnification.
@@
 We further visualize the gaps between model-merging and ensembling in fig.
-8(c).
+5(c).
 When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ is the peak for Cifar100).
@@
 A similar phenomenon is shown on different datasets, but with different peaks.
-Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab.
-3.
-From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent.
-It can also outperform logit ensembling, and cannot outperform feature ensembling.
-But generally it cannot work very well as on Cifar100.
-It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec.
-5.3, the model-merging can work to some extent as it can act similarly as regularization.
-It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
-Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$.
-This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect.
-However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
-Model # Perf Ave xM Perf Uni Soups Grd Soups Ens Lgt Ens Fts Ens Fts* Grd Ens Lgt Grd Ens Fts
-2 models 59.87 5.60 6.70 6.99 1.59 6.94 6.45 1.52 6.96
-3 models 59.85 5.44 7.15 7.05 0.78 7.38 6.78 0.78 7.21
-5 models 59.89 5.07 6.68 6.83 0.30 6.95 6.46 0.35 7.21
-7 models 59.90 4.80 6.77 6.80 0.24 6.73 6.26 0.21 6.74
-10 models 59.90 4.81 6.86 6.92 0.13 6.92 6.22 0.11 6.95
 7 Conclusion
@@
 We anticipate that these insights will stimulate further research into model-merging strategies and support the development of more interpretable and robust model-averaging methodologies.
-Supplementary Material
-8 Experimental Settings
+Appendix A More Figure Illustrations
+Figure 6: An illustration showing the magnitudes and variances of model outputs increasing over the increment of weights’ magnitudes and variances, making the prediction no longer fall into the correct side of the decision boundary.
+Figure 7: The examples of the first layer weights visualization of Resnet50 with nonlinearity inside the model.
+Appendix B Experimental Settings
 Table 2: The table shows dataset size, class number, and link for each of the dataset. In particular, ChestXRay dataset has 14 labels and each of them is a binary classification task.
-Datasets Sizes Class # Links
-Cifar10 60,000 10 https://www.cs.toronto.edu/~kriz/cifar.html
-Cifar100 60,000 100 https://www.cs.toronto.edu/~kriz/cifar.html
-PathMNIST 107,180 9 https://medmnist.com/
-DermaMNIST 10,015 7 https://medmnist.com/
-CelebA 202,599 2 https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
-TinyImageNet 100,000 200 https://huggingface.co/datasets/zh-plus/tiny-imagenet
-ChestXRay 112,120 14$\times$2 https://www.kaggle.com/paultimothymooney/chest-xray-pneumonia
 For the experimental settings of model weight patterns.
-We train one single linear layer on both Cifar100 [14] and Tiny ImageNet [15] datasets with momentum SGD as the optimizer.
+We train one single linear layer on both Cifar100 Krizhevsky (2009) and Tiny ImageNet Le and Yang (2015) datasets with momentum SGD as the optimizer.
 The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1.
@@
 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
-In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [35] for weight-averaged model-merging, which is the focus of our paper.
+In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups Wortsman et al.
+(2022) for weight-averaged model-merging, which is the focus of our paper.
 We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances.
-For the models considered, we have ResNet [9] (ResNet18 to ResNet152), DenseNet [10] (DenseNet121 to DenseNet201), VGG [24] (VGG11 to VGG19, including with and without batch normalization), ViT [26, 5] and DeiT [29].
-For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [39, 40], Celeba [17], Tiny ImageNet and ChestXRay [33, 18].
-The information of these datasets can be find in Tab.
-4.
+For the models considered, we have ResNet He et al.
+(2016) (ResNet18 to ResNet152), DenseNet Huang et al.
+(2017) (DenseNet121 to DenseNet201), VGG Simonyan (2014) (VGG11 to VGG19, including with and without batch normalization), ViT Steiner et al.
+(2021); Dosovitskiy (2020) and DeiT Touvron et al.
+(2021).
+For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST Yang et al.
+(2021, 2023b), Celeba Liu et al.
+(2015), Tiny ImageNet and ChestXRay Wang et al.
+(2017); Ma et al.
+(2019).
+The information of these datasets can be find in table 2.
 Specially, ChestXRay-14 is a chest X-ray imaging dataset designed for multi-label classification.
@@
 Also, we uniformly train 10 models and perform the model-merging and ensembling under exactly same conditions (learning rates, step strategies and optimizers, model numbers, etc).
-9 Theoretical Proofs
+Appendix C Theoretical Proofs of Weight Magnitude/Variance Expansion Effects Output Magnitude/Variance
 For Eq. 14 in the paper, we aim to prove that:
@@
 Thus, the inequality holds true.
-For the Property 1 and Theorem 1, according to [34], we have:
+For the Property 1 and Theorem 1, according to Wang et al. (2024), we have:
 Lemma 1: Given a fully-connected network, its weight matrix $\mathbf{W}^{(m)}$ satisfies, for any $\tau>0$,
@@
 This result follows by iteratively applying the inequalities, concluding the proof. ∎
-10 More Visualization of Model Weight Patterns
+Appendix D More Visualization of Model Weight Patterns
+D.0.1 Visualize Patterns Captured in Weights of 100 Classes for Cifar100 Dataset
 Figure 8: Average of all images within each class of Cifar100 dataset (all 100 classes).
@@
 We further visualize the linear classifiers after pair-wise merging between the first 50 classes and the last 50 classes in Cifar100 dataset as fig.
-11.
+10.
 We can see the combinations of two templates, e.g.
 “motorcycle+woman”.
+D.0.2 Visualization of Deep Convolutional Kernels Trained on Cifar100 Dataset
 Figure 11: The first layer weights visualization of ResNet50.
@@
 The patterns inside ResNet50 is more human-friendly to be observed than VGG19, as ResNet50 models have larger kernel sizes.
+D.0.3 Visualize Patterns Captured in Weights of 200 Classes for TinyImageNet Dataset
 Figure 13: Average of all images within each class of TinyImageNet dataset (the first 100 classes).
@@
 We can observe the correlations between averaged images and the learned patterns show a similar phenomenon.
-11 More Experimental Results of Averaging on Weights V.S. Averaging on Features
-Figure 18: The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100.
-Figure 19: The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
-Figure 20: The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100.
-Figure 21: The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+Appendix E More Experimental Results of Averaging on Weights V.S. Averaging on Features
+E.1 Merging Models Trained over Different Tasks
+Table 3: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods.
+Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
+E.g.
+Model 2 trained on Cifar10 receives 5.84% accuracy on Cifar100 dataset.
+Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
+Cifar10 60,000 5.84 49.63 69.93 69.93
+PathMNIST 107,180 1.29 24.25 69.88 70.00
+DermaMNIST 10,015 5.77 60.37 69.06 69.36
+Celeba 202,599 1.03 49.33 69.48 69.61
+TinyImageNet 100,000 18.16 61.02 58.40 58.52
+ChestXRay 112,120 15.18 61.38 68.18 68.33
+We also explore model-merging between different well-trained models on different tasks, as shown in table 3.
+Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling.
+Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy).
+It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features.
+Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling.
+Maybe because it is too small.
+In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along.
+It illustrates that the dataset size also matters.
+The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).
+E.2 2 Models, 3 Models, 5 Models and 7 Models for Model Merging/Ensembling Accuracy Comparison
+Figure 17: (a) The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+(b) The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+Figure 18: (a) The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100.
+(b) The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
 The 2 models, 3 models, 5 models and 7 models for model merging/ensembling accuracy comparison between different architectures on Cifar100 are shown via fig.
-18, Fig.
-19, Fig.
-20 and Fig.
-21, respectively.
+17(a), fig.
+17(b), fig.
+18(a) and fig.
+18(b), respectively.
 They all show that the models with ensembling on logits generally receive the best results and uniform soups receive the worst results.
@@
 ChestXRay 80.60 79.22 80.43 81.61 80.99 80.43 80.43
-Also, we show the dataset-wise exploration of 2 models, 3 models, 5 models and 7 models merging/ensembling in Tab.
-5, Tab.
-6, Tab.
-7 and Tab.
-8, respectively.
-The computation consumptions in Floating Point Operations (FLOPs) and Parameter Numbers are in Tab.
-9.
+Also, we show the dataset-wise exploration of 2 models, 3 models, 5 models and 7 models merging/ensembling in table 4, table 5, table 6 and table 7, respectively.
+The computation consumptions in Floating Point Operations (FLOPs) and Parameter Numbers are in table 8.
 Table 8: Comparison of FLOPs and Parameter Numbers for Various Models.
@@
 10 models 66.79 8.91 65.99 73.93 72.58 65.99 65.77
-We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab.
-10.
+We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in table 9.
 We can also see the more models are merged, the worse results from uniform soups can get.
@@
 Thus, it shows the ViT models are not suitable for model-merging in the tested scenarios.
-12 More Experimental Results of Model Magnitude Changes
-Figure 23: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar100 dataset.
-Figure 24: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar100 dataset.
-Figure 25: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar100 dataset.
-Figure 26: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
-The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
-Figure 27: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset.
-Figure 28: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar10 dataset.
-Figure 29: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar10 dataset.
-Figure 30: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar10 dataset.
-Figure 31: Accuracy vs. Magnification Factor across Methods with 10 models on Cifar10 dataset.
-Figure 32: Uniform soup - Logits Ensemble Accuracy vs. Magnification Factor across different model numbers on Cifar10 dataset.
+Appendix F More Experimental Results of Model Magnitude Changes
+F.1 Model-merging can mitigate weight magnitude issue on Tiny ImageNet
+Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab.
+10.
+From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent.
+It can also outperform logit ensembling, and cannot outperform feature ensembling.
+But generally it cannot work very well as on Cifar100.
+It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec.
+5.3, the model-merging can work to some extent as it can act similarly as regularization.
+It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
+Table 10: Model-merging for mitigate weight magnitude issue with different VGG19 model numbers on Tiny ImageNet with the magnitude factor ×$\times 100$ (accuracy is the metric).
+This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect.
+However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
+Model # Perf Ave xM Perf Uni Soups Grd Soups Ens Lgt Ens Fts Ens Fts* Grd Ens Lgt Grd Ens Fts
+2 models 59.87 5.60 6.70 6.99 1.59 6.94 6.45 1.52 6.96
+3 models 59.85 5.44 7.15 7.05 0.78 7.38 6.78 0.78 7.21
+5 models 59.89 5.07 6.68 6.83 0.30 6.95 6.46 0.35 7.21
+7 models 59.90 4.80 6.77 6.80 0.24 6.73 6.26 0.21 6.74
+10 models 59.90 4.81 6.86 6.92 0.13 6.92 6.22 0.11 6.95
+F.2 2 Models, 3 Models, 5 Models and 7 Models Accuracy V.S. Different Magnification Factors
+Figure 20: (a) Accuracy vs. Magnification Factor across Methods with 2 models on Cifar100 dataset. (b) Accuracy vs. Magnification Factor across Methods with 3 models on Cifar100 dataset.
+Figure 21: (a) Accuracy vs. Magnification Factor across Methods with 5 models on Cifar100 dataset. (b) Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
+The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via fig. 20(a), fig. 20(b), fig. 21(a) and fig. 21(b), respectively.
+Figure 22: (a) Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset. (b) Accuracy vs. Magnification Factor across Methods with 3 models on Cifar10 dataset.
+Figure 23: (a) Accuracy vs. Magnification Factor across Methods with 5 models on Cifar10 dataset. (b) Accuracy vs. Magnification Factor across Methods with 7 models on Cifar10 dataset.
+Figure 24: (a) Accuracy vs.
+Magnification Factor across Methods with 10 models on Cifar10 dataset.
+(b) Uniform soup - Logits Ensemble Accuracy vs.
+Magnification Factor across different model numbers on Cifar10 dataset.
 Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs.
 magnification factor on Cifar10 are shown via fig.
-27, Fig.
-28, Fig.
-29, Fig.
-30 and Fig.
-31, respectively.
+22(a), fig.
+22(b), fig.
+23(a), fig.
+23(b) and fig.
+24(a), respectively.
 Moreover, we also visualize the uniform soup and logits ensemble accuracy vs.
 different magnification factors across different model numbers on the Cifar10 dataset in fig.
-32.
+24(b).
 We can observe similar phenomenons as discussed in the apepr.